### v1

# Quantitative estimates: How well does the discrete Fourier transform approximate the Fourier transform on $\mathbb{R}$?

Abstract
In order to compute the Fourier transform of a function $f$ on the real line numerically, one samples $f$ on a grid and then takes the discrete Fourier transform. We derive exact error estimates for this procedure in terms of the decay and smoothness of $f$. The analysis provides a new recipe of how to relate the number of samples, the sampling interval, and the grid size.
keywords:
Discrete Fourier transform, FFT, approximation rate, weight functions, amalgam space.
{MSCcodes}
42QA38,65T05,94A12,43A15
1 Introduction
The fast Fourier transform (FFT) is widely used in the applied sciences for the numerical approximation of the Fourier transform
$\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}x\xi}\mathrm{d}x$ (1)
from sampled values of $f$. Despite the overwhelming success of this approximation, there are surprisingly few rigorous investigations and, in our view, still substantial theoretical gaps when it comes to error estimates. As the FFT is a fast algorithm that computes the discrete Fourier transform, we are actually asking: how well does the discrete Fourier transform approximate the Fourier transform [8]?
Engineers and numerical practitioners compute the Fourier transform of a function $f$ on the real line as follows:
  • (i)
    Sample $f$ on an interval of length $p$ at $n$ equispaced points $hj$, for $j\in\mathbb{Z}$, $-\frac{n}{2}<j\leq\frac{n}{2}$, with step size $h$.
  • (ii)
    It is then taken for granted that the scaled discrete Fourier transform of length $n$ provides an approximation for the values of the continuous Fourier transform of $f$ at $k/p$, i.e.,
    $\hat{f}\big{(}\tfrac{k}{p}\big{)}\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\,,\qquad k\in\mathbb{Z},\quad-\frac{n}{2}<k\leq\frac{n}{2}\,.$ (2)
This practice is usually unquestioned and works like a charm. It is motivated by following approximations with quadrature and truncation,
$\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}\xi x}\mathrm{d}x\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\,.$ (3)
For $\xi=\frac{k}{p}=\frac{k}{hn}$, the right-hand side is the scaled discrete Fourier transform (2). Yet one must ask what is the precise relation between the computed values $h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}$ and the samples $\hat{f}(\frac{k}{p})$ of the actual Fourier transform. Ideally one can provide reasonable error estimates for this approximation, and we formulate two questions:
  • (Q1)
    The numerical procedure depends on the number of samples $n$, the step size $h$, and the length $p$ of the sampled interval via the identity $p=hn$. How should $h,n$ and $p$ be chosen in dependence of the function class of $f$?
  • (Q2)
    What can be said about the asymptotic decay of the deviation in (2) when $n,p\rightarrow\infty$ and $h\to 0$?
State-of-the-art. To the best of our knowledge, there are surprisingly few investigations and even fewer quantitative error estimates in the literature.
The “Owner’s Manual for the Discrete Fourier Transform” [4, Section 6] contains a substantial chapter on the error of the discrete Fourier transform. There the error is quantified for rather special function classes, e.g., functions with compact support or functions on $\mathbb{R}$ with exponential decay with some type of finite order of smoothness conditions [4, Theorem 6.6] that are rarely met, see also our Section 3.5.
[…]
Since the right-hand side of (2) approximates the integral $\int_{-p/2}^{p/2}f(x)e^{-2\pi ikx/n}\mathrm{d}x$ by means of a Riemann sum or the trapezoidal rule, one should mention the recent error analyses [7, 11, 23, 25, 24] for the trapezoidal rule as relevant in this context. However, these error estimates require exponential decay and analyticity and do not even mention the Fourier transform.
In this paper we derive estimates for the error between the Fourier transform on the real line and the discrete Fourier transform for a large class of functions with polynomial or sub-exponential decay and finite smoothness. Our results confirm that, even under such very mild assumptions on the decay and smoothness, the folklore procedure works well and is successful. Our analysis also provides the new insight that the optimal spacing should depend on the decay of $f$ and $\hat{f}$. In view of (Q1), we identify an optimal relation between the step size $h$, the number of samples $n$, and the length of the sampling interval $p$ (the step size $\frac{1}{p}$ in frequency). To answer (Q2), we derive precise error estimates with explicit constants for function classes that are significantly larger than exponentially decaying, analytic functions. These error estimates are confirmed by numerical simulations utilizing the FFT which yield the precise asymptotics predicted by our theoretical estimates.
Our contribution. We proceed with a formal presentation of the numerical procedure and an explicit exposition of our results.
We first introduce the appropriate notation. For $n\in\mathbb{N}$, set
$[n]:=\{j\in\mathbb{Z}:-\tfrac{n}{2}<j\leq\tfrac{n}{2}\}\,.$
The discrete Fourier transform $\operatorname{\mathcal{F}}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ of $y=(y_{j})_{j\in[n]}\in\mathbb{C}^{n}$ is
$\operatorname{\mathcal{F}}\!y=\bigg{(}\frac{1}{\sqrt{n}}\sum_{j\in[n]}y_{j}e^{-2\pi\mathrm{i}\frac{kj}{n}}\bigg{)}_{k\in[n]}\,.$ (4)
Define the scaled sampling of a function $f$ with step size $h$ and length $n$ as
$f_{h,n}:=\Big{(}\sqrt{h}f(hj)\Big{)}_{j\in[n]}\in\mathbb{C}^{n}\,.$
Throughout the text we use the relation $p=hn$ between the interval length $p$, the step size $h$ and the number of samples $n$.
Our aim is to bound the approximation error
$E^{[n]}_{h}(f):=\biggl{(}\frac{1}{p}\sum_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}^{2}\bigg{)}^{1/2}=\big{\|}\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\big{\|}\,.$ (5)
The normalization of $f_{h,n}$ and the factor $\frac{1}{p}$ in the definition of the error may seem artificial, but they are justified and natural in view of the time-frequency symmetry of the error
$E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}})\,.$
See Appendix A for the proof. The symmetry between $f$ and $\hat{f}$ will also appear in the error estimates.
The following error estimate is a simple consequence of our main result for polynomial decay.
Theorem 1.1.
Let $a,b>1$. If $f$ and $\hat{f}$ are continuous and satisfy polynomial decay of the form [^1] We write $\lesssim$ if the left-hand-side is bounded by a constant times the right-hand-side. If $\lesssim$ and $\gtrsim$ both hold, then we write $\asymp$. The dependency of the constants on other parameters shall be clarified or is clear from the context. $|f(x)|\lesssim(1+|x|)^{-a}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$, then for all $\alpha<a-\frac{1}{2}$, $\beta<b-\frac{1}{2}$
$E^{[n]}_{h}(f)\lesssim p^{-\alpha}\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,,$
where the constants in $\lesssim$ are independent of $h$, $n$, $p$.
The error bound is balanced if $p^{-\alpha}\asymp h^{\beta}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$, and for this choice $E^{[n]}_{h}(f)\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}$. This is the optimal choice for the relation between $h$, $p$, and $n$, see Corollary 3.9. Our numerical simulations in Section 2 using the FFT suggest that the bounds in Theorem 1.1 are sharp.
For the approximation procedure to make sense, we need the Fourier transform to be defined pointwise and continuous, so that $f\in L^{1}(\mathbb{R})$ is a reasonable minimal assumption. One may argue that we also make certain assumptions on $\hat{f}$, which is the very object we want to compute. If one does not like the polynomial decay condition of $\hat{f}$ in Theorem 1.1, one can replace it by a stronger condition on smoothness of $f$. Precisely, if the derivatives $f^{(l)}$ are in $L^{1}(\mathbb{R})$ for $l=0,\dots,b$ with $b\in\mathbb{N}$, then it follows that $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$.
Sub-exponential decay in time and in frequency occurs importantly in the theory of test functions and distributions (under the name of Gelfand-Shilov spaces or ultra test functions) [3]. In that case, we offer the following version.
Theorem 1.2 (sub-exponential decay).
Let $r,s>0$ and $0<\alpha,\beta\leq 1$. If $f$ and $\hat{f}$ are continuous and satisfy the decay conditions $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim e^{-s|x|^{\beta}}$, then for all $r^{\prime}<r$,s<$s^{\prime}<s$,
$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}+e^{-s^{\prime}(2h)^{-\beta}}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|e^{s|\xi|^{\beta}}\,.$
For $r^{\prime}\to r$ and $s^{\prime}\to s$ the error is balanced if $e^{-r(\frac{p}{2})^{\alpha}}\asymp e^{-s(2h)^{-\beta}}$, which is satisfied for $h=n^{-\frac{\alpha}{\alpha+\beta}}\left(\tfrac{s}{r}\right)^{\frac{1}{\alpha+\beta}}2^{\frac{\alpha-\beta}{\alpha+\beta}}$. This condition on $h$ is the optimal choice for the relation between $h$, $p,$ and $n$, see Corollary 3.15.
We also provide bounds for mixed decay.
Theorem 1.3 (mixed decay).
Let $r>0$ and $0<\alpha\leq 1$. If $f$ and $\hat{f}$ are continuous and satisfy $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$,
$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,.$ (6)
For all practical purposes, the length of the sampled interval $p$ is sufficiently large, so that $e^{-r^{\prime}(\frac{p}{2})^{\alpha}}$ is negligible (and may for instance even be close to machine precision). Then the error is dominated by the polynomial term $h^{\beta}$.
In the main theoretical part in Section 3, we will derive more general results by quantifying the decay of a function by means of its membership in a Wiener amalgam space. This is the canonical function class when sampling is involved [9, 10, 12].
Software to plot the discrete Fourier transform commonly performs interpolation of the computed data vector $\operatorname{\mathcal{F}}\!f_{h,n}$. In order to simulate this process, we use three representative interpolation procedures and derive bounds on their deviation from $\hat{f}$ on the real line. One of them is based on the cardinal sine function $\operatorname*{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$ that is widely used in digital signal processing. The computation of $\operatorname{\mathcal{F}}\!f_{h,n}$ gives rise to the approximation of $\hat{f}$ on the real line by
$\hat{f}(\xi)\approx\sqrt{p}\sum_{k\in[n]}(\mathcal{F}f_{h,n})(k)\operatorname*{sinc}(p\xi-k)=h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\xi-k)\,.$
So the error is then $\|\hat{f}-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\cdot-k)\|_{L^{2}(\mathbb{R})}$. We will show in Section 5, Theorem 4.1, that this error obeys the same bounds as in Theorems 1.1, 1.2, and 1.3. The proof uses the estimates for $E^{[n]}_{h}(f)$.
Outline: Section 2 is dedicated to numerical experiments using the FFT and fully supports the theoretical results of Theorems 1.1 and 1.3. The full versions of the theoretical results on the approximation rates for spaces of polynomial and sub-exponential decay in time and in frequency are derived in Section 3. The approximation of $\hat{f}$ on the real line is discussed in Section 4. We investigate minimal requirements on $f$ and $\hat{f}$ for convergence in Section 5.
2 Numerical experiments for polynomial decay
Since our priority is the practical impact of the error estimates, we first offer numerical evidence that the error bounds of Theorem 1.1 are sharp for polynomial weights.
2.1 Polynomial decay in time and in frequency
In view of Theorem 1.1, we construct a family $f^{a,b}$ of functions with exact polynomial decay $a$ in time and $b$ in frequency such that each function and its Fourier transform can be numerically evaluated with sufficient accuracy.
We use linear combinations of shifts of the cardinal B-splines defined by $B_{1}:={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$, and
$B_{b+1}:=B_{b}*B_{1},\qquad b=1,2,\ldots\,.$
Each $B_{b}$ is a piecewise polynomial function of degree $b-1$ that is $b-2$-times continuously differentiable with support $[-\frac{b}{2},\frac{b}{2}]$, see [20] and also [19, Section 9.1].
Consider $(c_{k})_{k\in\mathbb{Z}}\subseteq\mathbb{C}$ given by its nonzero entries
$c_{0}=\frac{1}{2},\qquad c_{k}=\frac{1}{k\pi\mathrm{i}},\text{ for }k\in 2\mathbb{Z}+1\,,$
so that $c_{2k}=0$. We study the family of functions
$f^{a,b}(x):=\sum_{k\in\mathbb{Z}}c^{a}_{k}\,B_{b}(x-k),\qquad a,b=1,2,\ldots\,,$ (7)
which consists of locally finite sums. Therefore the point evaluations can be evaluated accurately in numerical experiments, and they satisfy $|f^{a,b}(x)|\lesssim(1+|x|)^{-a}$.
To compute the Fourier transform of $f^{a,b}$, we define the Fourier series $u_{a}(\xi):=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}$. As $\widehat{B_{b}}=\operatorname*{sinc}^{b}$, we observe
$\widehat{f^{a,b}}(\xi)$ $=\operatorname*{sinc}(\xi)^{b}\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}=\operatorname*{sinc}(\xi)^{b}u_{a}(\xi)\,.$
The estimate $|\operatorname*{sinc}^{b}(\xi)|\lesssim(1+|\xi|)^{-b}$ implies $|\widehat{f^{a,b}}(\xi)|\lesssim(1+|\xi|)^{-b}$, thus $f^{a,b}$ satisfies the assumptions of Theorem 1.1.
By a short computation we find $u_{1}(\xi)={\bf 1}_{(-\frac{1}{2},0)}(\xi)+\frac{1}{2}{\bf 1}_{\{-\frac{1}{2},0\}}(\xi)$ for $\xi\in[-\frac{1}{2},\frac{1}{2})$, and then $u_{a+1}$ is the cyclic convolution of $u_{a}$ and $u_{1}$, i.e.,
$u_{a+1}(\xi)=\int_{\mathbb{R}/\mathbb{Z}}u_{a}(\eta)u_{1}(\xi-\eta)\mathrm{d}\eta,\qquad a=1,2,\ldots\,.$
For $a=2,3$ we have $u_{2}(\xi)=|\xi|$ $u_{a}$ for $\xi\in[-\frac{1}{2},\frac{1}{2})$, and $u_{3}(\xi)=\operatorname*{sign}(\xi)\xi^{2}-\frac{1}{2}\xi+\frac{1}{8}$. Thus, we can evaluate the point evaluations of $f^{a,b}$ and $\widehat{f^{a,b}}$ exactly, both analytically and numerically, and the decay is labeled by $a$ and $b$. For the numerical simulations we compute the samples $\widehat{f^{a,b}}(k/p)$ and compare them with the discrete Fourier transform of a sampled version of $f^{a,b}$.
Since the error estimates of Theorem 1.1 work for every parameter $\alpha<a-1/2$ and $\beta<b-1/2$, we will plot the error curves with the values $\alpha=a-\frac{1}{2}$ and $\beta=b-\frac{1}{2}$ and compare our numerical observations with the error rate $n^{-\frac{\alpha\beta}{\alpha+\beta}}$ for $h=n^{-\frac{\alpha}{\alpha+\beta}}$. The infinitesimal difference between $\alpha<a-1/2$ and $\alpha=a-1/2$ is not visible.
To compute the discrete Fourier transform, we choose $n=2^{l}$, $l=10,\ldots,20$ in Theorem 1.1 and then apply the FFT to the samples $f^{a,b}_{h,n}$. The FFT is taken from the Julia package FFTW.jl with $64$-bit floating point arithmetics. The package provides a binding to the C library FFTW [13].
The approximation rates, predicted by Theorem 1.1, are $n^{-\frac{\alpha\beta}{\alpha+\beta}}\asymp h^{\beta}$, see Table 1 for some combinations of $a$ and $b$. To stick to $L^{1}(\mathbb{R})$ functions, we skip $a=1$.
$a$ $b$ $\alpha=a-\frac{1}{2}$ $\beta=b-\frac{1}{2}$ $h=n^{-\frac{\alpha}{\alpha+\beta}}$ $p=nh=n^{\frac{\beta}{\alpha+\beta}}$ $E^{[n]}_{h}(f^{a,b})\asymp h^{\beta}=n^{-\frac{\alpha\beta}{\alpha+\beta}}$
$2$ $2$ $3/2$ $3/2$ $n^{-1/2}$ $n^{1/2}$ $n^{-3/4}$
$2$ $3$ $3/2$ $5/2$ $n^{-3/8}$ $n^{5/8}$ $n^{-15/16}$
$2$ $4$ $3/2$ $7/2$ $n^{-3/10}$ $n^{7/10}$ $n^{-21/20}$
$3$ $2$ $5/2$ $3/2$ $n^{-5/8}$ $n^{3/8}$ $n^{-15/16}$
$3$ $3$ $5/2$ $5/2$ $n^{-1/2}$ $n^{1/2}$ $n^{-5/4}$
$3$ $4$ $5/2$ $7/2$ $n^{-5/12}$ $n^{7/12}$ $n^{-35/24}$
Table 1: List of the relevant values of Theorem 1.1 for $f^{a,b}$, so that $E^{[n]}_{h}(f^{a,b})$ is bounded by $h^{\beta}=p^{-\alpha}=n^{-\frac{\alpha\beta}{\alpha+\beta}}$.
Our numerical experiments as illustrated in Figure 1 align with the theoretical bounds in Theorem 1.1 and suggest that they are sharp.
Figure 1: Logarithmic plots of $E^{[n]}_{h}(f^{a,b})$ for various $a$ and $b$. Points are computed numerically for $n=2^{l}$, $l=10,\ldots,20$. Reference lines represent the theoretical slopes from Theorem 1.1 listed in Table 1.
Remark 2.1.
The closed form expressions for $u_{a}(\xi)=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}$ are key to our numerical results: Without a closed form expression of $u_{a}$ the series in (7) would need to be truncated for evaluation. The truncation error is accompanied by competing floating point errors due to the summation. We found it difficult to identify a sufficiently accurate truncation that yields reliable numerical estimates for $E^{[n]}_{h}(f^{a,b})$.
2.2 Polynomial decay in frequency
To test Theorem 1.3, we consider $f(x)=e^{-2\pi|x|}$. Its Fourier transform is $\hat{f}(\xi)=\frac{1}{\pi(1+\xi^{2})}$, so that $r=2\pi$, $\alpha=1$, and $|\widehat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b=2$.
For increasing values of $n$ in numerical experiments with the FFT, we choose successively $p=n^{1/2}$, $p=6n^{1/10}$, and $p=10$, so that $h=n^{-1/2}$, $h=6n^{-9/10}$, and $h=10n^{-1}$, respectively. Note that $p$ is sufficiently large, so that the error bound in Theorem 1.3 is expected to be governed by the polynomial term for a large range of $n$. Hence, we expect to observe the decay rate $h^{\frac{3}{2}}$, which is $n^{-\frac{3}{4}}$, $n^{-\frac{27}{10}}$, and $n^{-\frac{3}{2}}$, respectively. Figure 2 shows the results of our numerical experiments. Again, they match the theoretical prediction perfectly.
Figure 2: Logarithmic plots of $E^{[n]}_{h}(f)$ for $f(x)=e^{-2\pi|x|}$. Points are computed numerically for $n=2^{l}$, $l=7,\ldots,23$. We choose $p=n^{\frac{1}{2}}$, $p=6n^{1/10}$, and $p=10$. Reference lines represent the theoretical slope $h^{\beta}$ with $\beta=\frac{3}{2}$ that leads to $n^{-\frac{3}{4}}$, $n^{-\frac{27}{20}}$, and $n^{-\frac{3}{2}}$, respectively. Points align well with theoretical predictions.
3 Approximation rates
In this section we derive general and explicit bounds on the approximation error $E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|$ in (5) by decomposing the error into a time and a frequency component that are estimated separately. These components arise from approximating $\hat{f}(\tfrac{k}{p})$ first by a Riemannian sum and then truncating it as in (3),
$\hat{f}\big{(}\tfrac{k}{p}\big{)}\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}\approx h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=\sqrt{hn}(\operatorname{\mathcal{F}}\!f_{h,n})(k)\,.$ (8)
The left-hand side is the sampled (continuous) Fourier transform, the right-hand side is its approximation by the discrete Fourier transform. The idea is to estimate the discretization error $\hat{f}\big{(}\tfrac{k}{p}\big{)}-h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$ by the decay of $\hat{f}$, and the truncation $h\sum_{j\in\mathbb{Z}\setminus[n]}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$ by the decay of $f$.
3.1 Error decomposition into time and frequency components
The continuous functions on the real line are denoted by $\mathcal{C}=\mathcal{C}(\mathbb{R})$. We consider the Wiener amalgam space $W(\mathcal{C},\ell^{1})$ that consists of all continuous functions $f$ such that
$\|f\|_{W(\mathcal{C},\ell^{1})}:=\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|$ (9)
is finite, cf. [9, 10, 12]. If $f\in W(\mathcal{C},\ell^{1})$, then the sampling operator $f\mapsto(f(hj))_{j\in\mathbb{Z}}\in\ell^{1}(\mathbb{Z})$ is bounded for all $h>0$, and consequently the $p$-periodization
$\mathcal{P}\!_{p}f(x)=\sum_{l\in\mathbb{Z}}f(x+pl)$
converges absolutely and uniformly for every $p>0$. For $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, the Poisson summation formula $\sum_{l\in\mathbb{Z}}f(x+pl)=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi i\frac{k}{p}x}$ holds pointwise for all $x\in\mathbb{R}$, cf. [14, Lemma 4].
The following identity provides the foundation of our approach and will facilitate the error decomposition in a time and a frequency component.
Lemma 3.1.
If $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then, for $k\in[n]$,
$h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}=h\sum_{j\in[n]}(\mathcal{P}\!_{p}f)(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=(\mathcal{P}\!_{h^{-1}}\hat{f})(\tfrac{k}{p})\,.$ (10)
We may also write the second identity in terms of $(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}=\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}$.
Proof 3.2.
The first identity is the Poisson summation formula applied to the $h^{-1}$-periodization of $\hat{f}$. For the second identity, we partition $\mathbb{Z}$ into residue classes modulo $n$ and use $p=hn$. Then
$h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$ $=h\sum_{j\in[n]}\sum_{l\in\mathbb{Z}}f(hj+hnl)e^{-2\pi\mathrm{i}\frac{k}{hn}(hj+hnl)}$
$=h\sum_{j\in[n]}\sum_{l\in\mathbb{Z}}f(hj+pl)e^{-2\pi\mathrm{i}\frac{k}{hn}hj}$
$=h\sum_{j\in[n]}\mathcal{P}\!_{p}f(hj)e^{-2\pi\mathrm{i}\frac{k}{n}j}\,.$
Lemma 3.1 enables a reinterpretation of the two step approximation strategy via quadrature and truncation (8). The quadrature step compares the samples of the Fourier transform $\hat{f}(\frac{k}{p})$ to samples of the periodization of the Fourier transform $(\mathcal{P}\!_{h^{-1}}\hat{f})(\frac{k}{p})$, which coincide with the discrete Fourier transform of samples of the periodization of $f$, i.e., $\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}(k)$. The truncation compares the latter with the discrete Fourier transform $\operatorname{\mathcal{F}}\!f_{h,n}$. Symbolically,
$\hat{f}_{\frac{1}{p},n}\approx(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}=\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}\approx\operatorname{\mathcal{F}}\!f_{h,n}\,.$ (11)
More precisely, we split the pointwise error in two components,
$|\frac{1}{\sqrt{p}}\hat{f}(\tfrac{k}{p})-\operatorname{\mathcal{F}}\!f_{h,n}(k)|\leq\frac{1}{\sqrt{p}}|\hat{f}(\tfrac{k}{p})-\mathcal{P}\!_{h^{-1}}\hat{f}(\tfrac{k}{p})|+|\operatorname{\mathcal{F}}\Big{(}(\mathcal{P}\!_{p}f-f)_{h,n}\Big{)}(k)|\,.$ (12)
Taking norms and using the unitarity of $\operatorname{\mathcal{F}}$ we arrive at
$\begin{split}E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|\leq&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|\operatorname{\mathcal{F}}\Big{(}(\mathcal{P}\!_{p}f-f)_{h,n}\Big{)}\|\\ =&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(\mathcal{P}\!_{p}f-f)_{h,n}\|\,.\end{split}$ (13)
On the right hand side of this inequality we have (i) eliminated $\operatorname{\mathcal{F}}$ and (ii) split the error in a time component and a frequency component. Note that the symmetry of this decomposition matches the symmetry of the error $E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}})$, cf. Appendix A. To derive approximation rates, we must ensure that the samples of $f-\mathcal{P}\!_{p}f$ and $\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f}$ decay sufficiently fast. This can be accomplished by specifying decay conditions on $f$ and $\hat{f}$.
3.2 General weights
We quantify the decay of a function by the use of a weight function $v:\mathbb{R}\rightarrow(0,\infty)$. We consider weights $v$ that satisfy
$v\text{ is even: }v(-x)=v(x)\,,$ (14a)
$v\text{ is nondecreasing, i.e., }v(x)\leq v(y),\text{ for }|x|\leq|y|\,,$ (14b)
$v\text{ is submultiplicative, i.e., }v(x+y)\leq v(x)v(y)\text{ for }x,y\in\mathbb{R}\,,$ (14c)
$v\text{ is inverse square-summable, }\sum_{l\in\mathbb{Z}}{|v(l)|^{-2}}<\infty\,.$ (14d)
The standard examples of weights that satisfy conditions (14a) to (14d) are the polynomial weights $v_{\alpha}=(1+|x|)^{\alpha}$ for $\alpha>\frac{1}{2}$, and the (sub-)exponential weights $v_{r,\alpha}=e^{r|x|^{\alpha}}$, for $0<\alpha\leq 1$ and $r>0$.
Associated to $v$ are the weighted sequence spaces
$\ell^{q}_{v}:=\ell^{q}_{v}(\mathbb{Z}):=\{(c_{l})_{l\in\mathbb{Z}}\subseteq\mathbb{C}:\sum_{l\in\mathbb{Z}}|c_{l}|^{q}|v(l)|^{q}<\infty\}\,,$
and the weighted Lebesgue space $L^{2}_{v}:=L^{2}_{v}(\mathbb{R})$ of all functions $f$ such that $\|f\|_{L^{2}_{v}}=\|fv\|_{L^{2}}$ is finite. For a constant weight $v\equiv 1$ we omit the subscript and write $\ell^{p}$ and $L^{p}$.
The Wiener amalgam space $W(\mathcal{C},\ell^{q}_{v})$ consists of all continuous functions $f$ such that
$\|f\|_{W(\mathcal{C},\ell^{q}_{v})}:=\bigg{(}\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{q}|v(l)|^{q}\bigg{)}^{1/q}$ (15)
is finite, cf. [9, 10, 12].
If $v$ satisfies (14d) then $\ell^{2}_{v}\subseteq\ell^{1}$, and we derive the continuous embeddings $W(\mathcal{C},\ell^{2}_{v})\subseteq W(\mathcal{C},\ell^{1})\subseteq L^{1}(\mathbb{R})$.
In order to bound the right hand side of the error functional $E^{[n]}_{h}(f)$ in (13), we need to estimate the deviation of $f$ and $\hat{f}$ from their respective periodizations. We derive such bounds in terms of the function
$\Phi_{v}(p):=\bigg{(}2\sum_{m=0}^{\infty}|v(pm+\tfrac{p}{2})|^{-2}\bigg{)}^{1/2}\,.$ (16)
Observe that $\lim_{p\to\infty}\Phi_{v}(p)=0$ always holds for a weight $v$ satisfying (3.2). For polynomial and sub-exponential weights we obtain quantitative bounds on the decay of $\Phi_{v}(p)$ in Lemma 3.7 and in Lemma 3.13.
Lemma 3.3.
If $f\in W(\mathcal{C},\ell^{2}_{v})$, and $v$ satisfies (3.2) then
$\|f-\mathcal{P}\!_{p}f\|_{L^{2}(-\frac{p}{2},\frac{p}{2})}$ $\leq v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})},$ (i)
$\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ $\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$ (ii)
Part (i) is used only in Section 4 on interpolation, but Part (ii) is crucial for bounds on $E^{[n]}_{h}$.
Proof 3.4.
(i) We bound $(f-\mathcal{P}\!_{p}f)(x)$ by the Cauchy-Schwartz inequality,
$\bigg{|}\sum_{0\neq l\in\mathbb{Z}}f(x+pl)\bigg{|}^{2}\leq\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}\sum_{0\neq l\in\mathbb{Z}}\left|f(x+pl)\right|^{2}|v(x+pl)|^{2}\,.$ (17)
Since $v$ is even (14a) and nondecreasing (14b), for $|x|\leq\frac{p}{2}$, the first sum can be majorized by
$\begin{split}\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}&\leq\sum_{0\neq m\in\mathbb{Z}}|v(p|m|-\tfrac{p}{2})|^{-2}\\ &\leq 2\sum_{m=1}^{\infty}|v(pm-\tfrac{p}{2})|^{-2}=|\Phi_{v}(p)|^{2}\,.\end{split}$ (18)
Integration of (17) yields
$\int_{-\frac{p}{2}}^{\frac{p}{2}}\bigg{|}\sum_{0\neq l\in\mathbb{Z}}f(x+pl)\bigg{|}^{2}\mathrm{d}x$ $\leq|\Phi_{v}(p)|^{2}\int_{-\frac{p}{2}}^{\frac{p}{2}}\sum_{0\neq l\in\mathbb{Z}}\left|f(x+pl)\right|^{2}|v(x+pl)|^{2}\mathrm{d}x$
$=|\Phi_{v}(p)|^{2}\int_{|x|>\frac{p}{2}}\left|f(x)\right|^{2}|v(x)|^{2}\mathrm{d}x$
$\leq|\Phi_{v}(p)|^{2}\|f\|^{2}_{L^{2}_{v}}\,.$
Direct computations lead to $\|f\|_{L^{2}_{v}}\leq v(1)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}$.
(ii) We use (17) and (18) for $x=hj$ and sum over $j\in[n]$, so that $p=nh$ yields
$\sum_{j\in[n]}\bigg{(}\sum_{0\neq l\in\mathbb{Z}}\left|f(hj+pl)\right|\bigg{)}^{2}$ $\leq|\Phi_{v}(p)|^{2}\sum_{j\in[n]}\sum_{0\neq l\in\mathbb{Z}}\left|f(hj+pl)\right|^{2}|v(hj+pl)|^{2}$
$=|\Phi_{v}(p)|^{2}\sum_{j\in\mathbb{Z}\setminus[n]}\left|f(hj)\right|^{2}|v(hj)|^{2}$
$\leq|\Phi_{v}(p)|^{2}\sum_{j\in\mathbb{Z}}\left|f(hj)\right|^{2}|v(hj)|^{2}\,.$
[…]
$\sum_{j\in\mathbb{Z}}\left|f(hj)\right|^{2}|v(hj)|^{2}$ $\leq\lceil h^{-1}\rceil\sum_{l\in\mathbb{Z}}\sup_{x\in[l,l+1]}\left|f(x)\right|^{2}|v(|l|+1)|^{2}$
$\leq\lceil h^{-1}\rceil|v(1)|^{2}\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{2}|v(l)|^{2}\,.$
The observation $h\lceil h^{-1}\rceil\leq(1+h)$ concludes the proof.
After these preparations we can now formulate our main result. This is an error estimate for the approximation of the Fourier transform by the discrete Fourier transform with very general conditions on the decay of the function and its Fourier transform.
Theorem 3.5.
Let $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$. If the weights $v,w$ satisfy (3.2), then
$E^{[n]}_{h}(f)\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+w(1)\Phi_{w}(h^{-1})(1+\tfrac{1}{p})^{\frac{1}{2}}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (19)
Note that for $0<h\leq 1\leq p$ the first term does not depend on $h$ and the second term not on $p$. Therefore
$E^{[n]}_{h}(f)\leq 2^{\frac{1}{2}}v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+2^{\frac{1}{2}}w(1)\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (20)
We now prove Theorem 3.5.
Proof 3.6.
We apply Lemma 3.3 (ii) to both $f$ and $\hat{f}$. Recall the bound $E^{[n]}_{h}(f)\leq\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ in (13). Lemma 3.3 estimates $\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ and yields the first term of the error in (19). Applying Lemma 3.3 to $\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|$ by interchanging the roles of $f$, $p$, and $h$ with $\hat{f}$, $h^{-1}$, and $p^{-1}$, yields the second term. This concludes the proof.
Theorem 3.5 asserts a bound on the error $E^{[n]}_{h}(f)$ in terms of $\Phi_{v}(p)$ and $\Phi_{w}(h^{-1})$. For polynomial and sub-exponential weights the error estimates can be made much more explicit.
3.3 Polynomial weights
In this section we consider the polynomial weights $v_{\alpha}(x)=(1+|x|)^{\alpha}$. To bound $\Phi_{v_{\alpha}}(p)$ in (16), we make use of the Hurwitz zeta function $\zeta(s,t)=\sum_{m=0}^{\infty}(m+t)^{-s}$.
Lemma 3.7.
If $\alpha>\frac{1}{2}$, then
$\Phi_{v_{\alpha}}(p)\leq p^{-\alpha}\sqrt{2\zeta(2\alpha,\tfrac{1}{2})}\leq p^{-\alpha}2^{\alpha+1/2}\Big{(}1+\frac{1}{4\alpha-2}\Big{)}^{1/2}\lesssim p^{-\alpha}\,.$
Proof 3.8.
We factor out $p$ to obtain
$|\Phi_{v_{\alpha}}(p)|^{2}$ $=2\sum_{m=0}^{\infty}\left(1+\left|pm+\tfrac{p}{2}\right|\right)^{-2\alpha}\leq p^{-2\alpha}2\sum_{m=0}^{\infty}\left(\tfrac{1}{p}+m+\tfrac{1}{2}\right)^{-2\alpha}$
$\leq p^{-2\alpha}2\sum_{m=0}^{\infty}\left(m+\tfrac{1}{2}\right)^{-2\alpha}=p^{-2\alpha}2\zeta(2\alpha,\tfrac{1}{2})\,.$
The integral test for convergence of series provides, for $\alpha>1/2$,
$\zeta(2\alpha,\tfrac{1}{2})\leq 2^{2\alpha}+\int_{0}^{\infty}(x+\tfrac{1}{2})^{-2\alpha}\mathrm{d}x=2^{2\alpha}(1+\tfrac{1}{4\alpha-2})\,.$
For polynomial weights Theorem 3.5 takes the following shape.
Corollary 3.9 (polynomial weights).
Assume that $f\in W(\mathcal{C},\ell^{2}_{v_{\alpha}})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{v_{\beta}})$ for $\alpha,\beta>\frac{1}{2}$, and $0<h\leq 1\leq p$.
  • (i)
    Set $c_{s}=2^{2s+1}\left(1+\tfrac{1}{4s-2}\right)^{\frac{1}{2}}$. Then
    $E^{[n]}_{h}(f)\leq c_{\alpha}p^{-\alpha}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}+c_{\beta}h^{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\,.$
  • (ii)
    If the step size satisfies $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$, then
    $E^{[n]}_{h}(f)\lesssim h^{\beta}\big{(}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\big{)}\,.$
    Expressed with the number of samples, the error is $E^{[n]}_{h}(f)\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}$.
Proof 3.10.
(i) The assumption $h\leq 1\leq p$ implies that $\sqrt{1+h}\leq\sqrt{2}$ and $\sqrt{1+p^{-1}}\leq\sqrt{2}$. By definition $v_{\alpha}(1)=2^{\alpha}$. Using Lemma 3.7, the constant $v_{\alpha}(1)\Phi_{v_{\alpha}}(p)\sqrt{1+h}$ in Theorem 3.5 is then bounded by $2^{2\alpha+1}(1+1/(4\alpha-2))^{1/2}$. Likewise for the constant involving $w=v_{\beta}$. The error estimate now follows from Theorem 3.5.
(ii) We balance the terms $h^{\beta}$ and $p^{-\alpha}$. Since $h=\frac{p}{n}$, the choice $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$ leads to $h^{\beta}\asymp p^{-\alpha}$ and the overall error is of order $h^{\beta}$.
Corollary 3.9 answers the questions (Q1) and (Q2) raised in the introduction whenever $f$ decays polynomially in time and in frequency.
Remark 3.11.
In the literature, e.g. [16], one can find the choice $h\asymp\frac{1}{\sqrt{n}}$. It is optimal only for $\alpha=\beta$. If $\alpha\neq\beta$, then $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$ leads to better bounds. Our analysis reveals that the spacing should depend on the decay of $f$ and $\hat{f}$, and Corollary 3.9 explains how to choose the optimal parameters.
In our experience, amalgam spaces are the natural theoretical framework when dealing with sampling and periodization. Yet one may wish to have conditions that are easier to check in practice. One such condition is a pure decay condition $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ as in Theorem 1.1.
Proof 3.12 (Proof of Theorem 1.1).
Assume that $|f(x)|\lesssim(1+|x|)^{-a}$ and $\alpha<a-\frac{1}{2}$. We show that $f\in W(\mathcal{C},\ell^{2}_{v_{\alpha}})$. Hölder’s inequality yields
$\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{2}(1+|l|)^{2\alpha}\lesssim\sup_{x\in\mathbb{R}}|f(x)|^{2}(1+|x|)^{2a}\sum_{l\in\mathbb{Z}}(1+|l|)^{-2(a-\alpha)}\,.$ (21)
Since $a-\alpha>\frac{1}{2}$, the series $\sum_{l\in\mathbb{Z}}(1+|l|)^{-2(a-\alpha)}<\infty$ converges, so that
$\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}\lesssim\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a},$
and likewise for $\hat{f}$. Therefore Theorem 1.1 is a special case of Theorem 3.9.
3.4 Sub-exponential weights
Next, we consider sub-exponential weights of the form $v_{r,\alpha}(x)=e^{r|x|^{\alpha}}$ as defined in Section 3.2. We first find a bound for $\Phi_{v_{r,\alpha}}(p)$ in (16). We directly derive
$|\Phi_{v_{r,\alpha}}(p)|^{2}=2\sum_{m=0}^{\infty}e^{-2r\left(pm+\frac{p}{2}\right)^{\alpha}}=e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,.$ (22)
For fixed $0<\alpha\leq 1$, the series $\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}$ converges for every $p>0$. As a function of $p$, it is monotonically decreasing, so that
$|\Phi_{v_{r,\alpha}}(p)|^{2}\leq e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{1}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,,$ (23)
for $p\geq 1$. To derive more explicit bounds, we use the incomplete Gamma function $\Gamma(s,u)=\int_{u}^{\infty}t^{s-1}e^{-t}\mathrm{d}t$.
Lemma 3.13.
Assume that $0<\alpha\leq 1$ and $r>0$.
  • (i)
    If $p\geq 1$, then
    $\Phi_{v_{r,\alpha}}(p)\leq e^{-r(\frac{p}{2})^{\alpha}}\Big{(}2+\frac{1}{\alpha}\frac{e^{r2^{1-\alpha}}}{(r2^{1-\alpha})^{1/\alpha}}\Gamma(\tfrac{1}{\alpha},r2^{1-\alpha})\Big{)}^{1/2}\,.$
  • (ii)
    If $p\geq\big{(}\frac{2^{\alpha}}{2r\alpha}\big{)}^{1/\alpha}$, then
    $\Phi_{v_{r,\alpha}}(p)\leq e^{-r(\frac{p}{2})^{\alpha}}\Big{(}2+\frac{2^{\alpha}}{2\alpha^{2}rp^{\alpha}}\Big{)}^{1/2}\leq e^{-r(\frac{p}{2})^{\alpha}}\Big{(}2+\frac{1}{\alpha}\Big{)}^{1/2}\,.$
In particular, the factor $\big{(}2+\frac{2^{\alpha}}{2\alpha^{2}rp^{\alpha}}\big{)}^{1/2}$ tends to $\sqrt{2}$ when $p\rightarrow\infty$.
Proof 3.14.
According to (22), we must bound the series $\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}$. To verify Part (ii), consider $p\geq\big{(}\frac{2^{\alpha}}{2r\alpha}\big{)}^{1/\alpha}$. For $u:={2r(\frac{p}{2})^{\alpha}}$, the integral test for convergence of series yields
$\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}$ $\leq 1+\int_{0}^{\infty}e^{-u\left((2x+1)^{\alpha}-1\right)}\mathrm{d}x=1+\frac{1}{2}e^{u}\int_{1}^{\infty}e^{-ux^{\alpha}}\mathrm{d}x\,.$
A direct check reveals that $-\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},x^{\alpha}u)$ is a primitive of $x\mapsto e^{-ux^{\alpha}}$ and therefore $\int_{1}^{\infty}e^{-ux^{\alpha}}\,\mathrm{d}x=-\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},u)$. Hence, we derive
$\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}\leq 1+\frac{1}{2\alpha}\frac{e^{u}}{u^{1/\alpha}}\Gamma(\tfrac{1}{\alpha},u)\,.$ (24)
The condition $p\geq\big{(}\frac{2^{\alpha}}{2r\alpha}\big{)}^{1/\alpha}$ implies $u=2r(\frac{p}{2})^{\alpha}\geq\frac{1}{\alpha}\geq 1$. In that case, the incomplete Gamma function is bounded by
$\Gamma(\tfrac{1}{\alpha},u)\leq\frac{1}{\alpha}u^{\frac{1}{\alpha}-1}e^{-u}\,,$ (25)
see [18, Proposition 2.7 and Page 1275]. For completeness we reproduce the elementary argument:
Set $a=\alpha^{-1}$. The function $h(t)=(a-1)\ln t-t$ is strictly concave, so it is majorized by its tangent $h_{u}(t)=h(u)+h^{\prime}(u)(u-x)$. Therefore
[…]
where the last inequality is due to $(1-(a-1)/u)^{-1}\leq a$ for $u\geq a\geq 1$. This yields (25).
We substitute (25) into (24), use the assumption $u\leq 1/\alpha$, and obtain
$\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}$ $\leq 1+\frac{1}{2\alpha^{2}u}\leq 1+\frac{1}{2\alpha}\,.$
Inserting $u={2r(\frac{p}{2})^{\alpha}}$ and the assumption $p^{\alpha}\geq\frac{2^{\alpha}}{2r\alpha}$ lead to
$\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\leq 1+\frac{2^{\alpha}}{4\alpha^{2}rp^{\alpha}}\leq 1+\frac{1}{2\alpha}\,.$
(i) The bound for $p\geq 1$ is simpler and follows directly from (23) with the choice $u={2r(\frac{1}{2})^{\alpha}}$ and (24).
For sub-exponential weights Theorem 3.5 takes the following shape.
Corollary 3.15 (sub-exponential weights).
Assume that $0<\alpha,\beta\leq 1$, that $r,s>0$, and $0<h\leq 1\leq p$. If $f\in W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{v_{s,\beta}})$, then the following error estimates hold.
  • (i)
    There are constants $c_{r,\alpha},c_{s,\beta}>0$ such that
    $E^{[n]}_{h}(f)\leq c_{r,\alpha}e^{-r(\frac{p}{2})^{\alpha}}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})}+c_{s,\beta}e^{-s(2h)^{-\beta}}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{s,\beta}})}\,.$
    If $p\geq\big{(}\frac{2^{\alpha}}{2r\alpha}\big{)}^{1/\alpha}$, then we may choose $c_{r,\alpha}=e^{r}(4+\frac{2}{\alpha})^{1/2}$, and if $h\leq\big{(}\frac{2s\beta}{2^{\beta}}\big{)}^{1/\beta}$, then $c_{s,\beta}=e^{s}(4+\frac{2}{\beta})^{1/2}$.
  • (ii)
    If the step size satisfies $h=n^{-\frac{\alpha}{\alpha+\beta}}\left(\tfrac{s}{r}\right)^{\frac{1}{\alpha+\beta}}2^{\frac{\alpha-\beta}{\alpha+\beta}}$, then
    $E^{[n]}_{h}(f)\lesssim e^{-s(2h)^{-\beta}}\big{(}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{s,\beta}})}\big{)}\,.$
Proof 3.16.
The proof is analogous to the one for polynomial weights (Corollary 3.9). We only check the constants. According to Theorem 3.5 and Lemma 3.13, the constant $c_{r,\alpha}$ can be chosen as
$c_{r,\alpha}=v_{r,\alpha}(1)(2+\tfrac{1}{\alpha})^{1/2}\sqrt{1+h}\leq e^{r}(4+\tfrac{2}{\alpha})^{1/2}\,,$
where we used $\sqrt{1+h}\leq\sqrt{2}$. Likewise, we get the specification of $c_{s,\beta}$.
Proof 3.17 (Proof of Theorem 1.2).
Assume that $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $r^{\prime}=r-\epsilon<r$. Then $f\in W(\mathcal{C},\ell^{1}_{r-\epsilon,\alpha})$ for every $\epsilon>0$ with the same argument as in (21). Therefore the theorem is a direct consequence of Corollary 3.15 and this embedding.
3.5 Mixed weights
We next discuss sub-exponential decay of $f$ and polynomial decay of $\hat{f}$. Results for sub-exponential decay of $\hat{f}$ and polynomial decay of $f$ are obtained by switching the roles of $f$ and $\hat{f}$ in the results below.
Assume that $f\in W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{v_{\beta}})$. We feed the estimates of Lemma 3.7 and 3.13 into Theorem 3.5 and obtain
$E^{[n]}_{h}(f)\leq e^{r}\big{(}4+\tfrac{2}{\alpha}\big{)}^{1/2}e^{-r(\frac{p}{2})^{\alpha}}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})}+2^{2\beta+1}\big{(}1+\tfrac{1}{4\beta-2}\big{)}^{\frac{1}{2}}h^{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}$ (26)
provided that $p\geq\big{(}\frac{2^{\alpha}}{2r\alpha}\big{)}^{1/\alpha}$. Both terms contribute in a balanced manner if $e^{-r(\frac{p}{2})^{\alpha}}=h^{\beta}$. Let $\mathcal{W}$ be the Lambert $W$-function, i.e., the inverse of $t\mapsto te^{t}$. Using $p=nh$ we express $h$ as
$h=n^{-1}2(\tfrac{\beta}{\alpha r})^{1/\alpha}\mathcal{W}^{1/\alpha}(\tfrac{\alpha n^{\alpha}r}{2^{\alpha}\beta})\,.$
As for direct decay conditions, if $f$ and $\hat{f}$ are continuous and satisfy $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$,
$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,.$ (27)
This yields Theorem 1.3 in the introduction.
For the exponential case $\alpha=1$, a related bound for the sup-norm of $\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}$ is derived by Briggs and Henson [4, Theorem 6.6]. However, they impose extremely restrictive regularity requirements. For a comparable bound the periodization $\mathcal{P}_{p}(f{\bf 1}_{[-\frac{p}{2},\frac{p}{2})]})$ must be $b-2$-times continuously differentiable on $\mathbb{R}$. This holds only for those interval lengths $p$, where the derivatives at ±$\pm\frac{p}{2}$ coincide.
4 Interpolation
So far, we have approximated the Fourier samples $\hat{f}(\tfrac{k}{p})$ by the discrete Fourier transform $h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}kj/n}=\sqrt{p}\,(\mathcal{F}f_{h,n})(k)$, for $k\in[n]$. In this section we change the point of view: we now interpolate the vector $\mathcal{F}f_{h,n}$ and study how the interpolating function approximates $\hat{f}$ on whole real line $\mathbb{R}$. To do so, we use three exemplary interpolation schemes that are often used in approximation theory.
The standard procedure starts with a cardinal interpolating function: this is a nice function $\phi$ on $\mathbb{R}$ that satisfies the interpolation condition $\phi(k)=\delta_{k}$. The approximation of $\hat{f}$ on $\mathbb{R}$ from the discrete Fourier transform is then given by
$\Phi_{h,n}(\xi)=\sqrt{p}\,\sum_{k\in[n]}\mathcal{F}f_{h,n}(k)\phi(p\xi-k)=h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\phi(p\xi-k)\,.$ (28)
This expression is an approximation of $\hat{f}$ on $\mathbb{R}$ that satisfies $\Phi_{h,n}(\tfrac{k}{n})=\sqrt{p}\,\mathcal{F}f_{h,n}(k)$, and we seek to estimate the global error $\|\hat{f}-\Phi_{h,n}\|_{2}$.
As the cardinal interpolating function $\phi$ we will use the first two B-splines $B_{1}={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$ and $B_{2}(x)=(1-|x|){\bf 1}_{[-1,1]}(x)$ and the cardinal sine function $\operatorname*{sinc}(x)=\tfrac{\sin\pi x}{\pi x}$. Clearly, these functions satisfy $\phi(k)=\delta_{k}$. To keep the notion compact, we set $B_{0}:=\operatorname*{sinc}$. For $\phi=B_{1}$, then right-hand side of (28) is a step function, for $\phi=B_{2}$ we obtain a continuous, piecewise linear function (this is how many software would plot the approximating vector $\mathcal{F}f_{h,n}$), and for $\phi=\operatorname*{sinc}$ we obtain a smooth function whose Fourier transform has support in $[-1/2,1/2]$, which is commonly referred to as a bandlimited function.
4.1 Convergence rates
We first derive an error bound for the bandlimited approximation (28) with $\Phi=\operatorname*{sinc}$ for general weights. This is then specified to the polynomial and the sub-exponential weights, resulting in explicit error rates. The approximation rates for the $B$-splines (28) with $i=1,2$ are restricted to $\alpha\leq 1$ and $\alpha\leq 2$, respectively.
Theorem 4.1.
Assume that $0<h\leq 1\leq p$ and that the weights $v,w$ satisfy (3.2). If $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$, then
$\Big{\|}\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (29)
If $\operatorname*{sinc}$ is replaced with $B_{i}$, for $i=1,2$, then the estimate (29) bounds holds for the polynomial weights $v=v_{\alpha}$ subject to the restriction $1/2<\alpha\leq i$.
We note that the error is of the same form as in Theorem 3.5. In particular, for polynomial weights $v=v_{\alpha}$ and $w=v_{\beta}$ with $\alpha,\beta>\frac{1}{2}$ we obtain $\Phi_{v}(p)\lesssim p^{-\alpha}$ and $\Phi_{w}(h^{-1})\lesssim h^{\beta}$ (see Lemma 3.7). For sub-exponential weights $v=v_{r,\alpha}$ and $w=v_{s,\beta}$ with $0<\alpha,\beta\leq 1$ and $r,s>0$ we obtain $\Phi_{v}(p)\lesssim e^{-r(\frac{p}{2})^{\alpha}}$ and $\Phi_{w}(h^{-1})\lesssim e^{-s(\frac{1}{2h})^{\beta}}$ (see Lemma 3.13).
Proof 4.2.
The proof is based on the following three approximation steps:
$\hat{f}\,\overset{(s1)}{\approx}\,\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)$ $\,\overset{(s2)}{\approx}\,\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)$
$\,\overset{(s3)}{\approx}\,h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\cdot-k)\,.$ (30)
We derive bounds for the steps (s2) and (s3) with respect to general weights simultaneously for $i=0,1,2$.
(s3) For $i=0,1$, the system $\{\sqrt{p}B_{i}(p\cdot-k)\}_{k\in\mathbb{Z}}$ is orthonormal in $L^{2}(\mathbb{R})$ and it is still a Riesz sequence for $i=2$ with Riesz bounds independent of $p$ (this can be verified by direct computation). Therefore, we conclude that
$\Big{\|}\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\cdot-k)\Big{\|}_{L^{2}}$
$=\|\sum_{k\in[n]}\frac{1}{\sqrt{p}}\Big{(}\hat{f}(\frac{k}{p})-\mathcal{F}f_{h,n}(k)\Big{)}\,\sqrt{p}\,B_{i}(p\cdot-k)\|_{L}^{2}\asymp\big{\|}\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\big{\|}\ =E^{[n]}_{h}(f)\,,$
and this is just the discrete approximation error that has been treated in Theorem 3.5.
(s2) Using orthonormality or the Riesz basis property again, we observe
$\Big{\|}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}^{2}_{L^{2}}\asymp p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}\,.$
The monotonicity (14b) of the weight $w$ implies
$p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}\lesssim p^{-1}|w(\tfrac{n}{2p})|^{-2}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}|w(\tfrac{k}{p})|^{2}\,.$
The sum on the right-hand side is bounded as in estimates at the end of the proof of Lemma 3.3. We deduce
$p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}|w(\tfrac{k}{p})|^{2}\leq p^{-1}\lceil p\rceil|w(1)|^{2}\|\hat{f}\|^{2}_{W(\mathcal{C},\ell^{2}_{w})}\,.$
As $|w(\tfrac{n}{2p})|^{-1}=|w(\tfrac{1}{2h})|^{-1}\leq\big{(}2\sum_{m=0}^{\infty}|w(\frac{m}{h}+\frac{1}{2h})|^{-2}\big{)}^{1/2}=\Phi_{w}(h^{-1})$ we obtain
$\Big{\|}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
(s1) The error $\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\|_{L^{2}}$ relates to classical interpolation of $\hat{f}$ from its samples [5, 6]. Since we consider general weights only for $B_{0}=\operatorname*{sinc}$, we carry out the calculations for $B_{0}$. Direct computations and Plancherel’s formula lead to
$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\cdot-k)\Big{\|}^{2}_{L^{2}}$ $=\Big{\|}f-\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}\,\cdot\,\frac{k}{p}}{\bf 1}_{[-\frac{p}{2},\frac{p}{2}]}\Big{\|}^{2}_{L^{2}}$
$=\int_{|x|>\frac{p}{2}}\!|f(x)|^{2}\mathrm{d}x+\!\int_{-\frac{p}{2}}^{\frac{p}{2}}\!\bigg{|}f(x)-\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}x\frac{k}{p}}\bigg{|}^{2}\mathrm{d}x\,.$
The first term can be estimated as
$\int_{|x|>\frac{p}{2}}\left|f(x)\right|^{2}\mathrm{d}x\leq|v(\tfrac{p}{2})|^{-2}\|f\|^{2}_{L^{2}_{v}}\lesssim\Phi^{2}_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,,$
where we used the continuous embedding $W(\mathcal{C},\ell^{2}_{v})\subseteq L^{2}_{v}$ and $|v(\tfrac{p}{2})|^{-2}\leq\Phi^{2}_{v}(p)$.
For the second term we use the Poisson summation formula and apply Lemma 3.3(i), so that
$\int_{-\frac{p}{2}}^{\frac{p}{2}}\Big{|}f(x)-p^{-1}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}x\frac{k}{p}}\Big{|}^{2}\mathrm{d}x$ $=\|f-\mathcal{P}\!_{p}f\|^{2}_{L^{2}(-\frac{p}{2},\frac{p}{2})}\lesssim\Phi^{2}_{v}(p)\|f\|^{2}_{W(\mathcal{C},\ell^{2}_{v})}\,.$
This provides the required bound
$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$
To treat the case $i=1,2$ of piecewise constant and piecewise linear interpolation for the polynomial weights $v=v_{\alpha}$ and $w=v_{\beta}$, we use results from [15, Thm 3 and Lemma 3]. For $\alpha\leq i,i=1,2,$ these imply
$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim p^{-\alpha}\|f\|_{L^{2}_{v_{\alpha}}}\,.$ (31)
The continuous embedding $W(\mathcal{C},\ell^{2}_{v_{i}})\subseteq L^{2}_{v_{i}}$ yields $\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\|_{L^{2}}\lesssim p^{-i}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{i}})}$.
The final error is the sum of the three errors (s1), (s2), and (s3).
Remark 4.3.
The restriction to $\alpha=i$, for $i=1,2$, in Part (b) of Theorem 4.1 stems from the bound (31). For stronger polynomial decay $\alpha$, the splines $B_{1}$ and $B_{2}$ need to be replaced by a suitable function $\phi_{\alpha}$, see [5] for examples. We need at least that $\phi_{\alpha}$ interpolates on $\mathbb{Z}$ and that its integer shifts form a Riesz-sequence in $L^{2}$.
The proof for the interpolating function $\phi=\operatorname*{sinc}$ works for all polynomials weights $v_{\alpha}$ with $\alpha>\frac{1}{2}$. Although the $\operatorname*{sinc}$-function has a bad reputation in numerical analysis due to its poor decay, it works well in the context of Fourier approximation. In particular, since we use only finite sums, the slow decay of $\operatorname*{sinc}$ does not cause any convergence issues. We refer to [21, 22] for $\operatorname*{sinc}$-methods in numerical analysis.
5 Minimal requirements
In this section we investigate under which conditions the error $E^{[n]}_{h}(f)$ tends to zero. Although this question may not be of practical value, it is of fundamental interest to know for which function class the Fourier transform can be successfully approximated by the discrete Fourier transform. We answer this question for the norm $\sqrt{p}\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}$ representing the pointwise error.
Proposition 5.1.
If $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then
$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\,\sup_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}=0\,.$
Proof 5.2.
Following the general approach outlined in Section 3.1 we decompose the error $\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}$ into a time and a frequency component. As in (13), we have
$\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}\leq\|\hat{f}_{\frac{1}{p},n}-(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\|\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}-\operatorname{\mathcal{F}}\!f_{h,n}\|_{\infty}\,.$ (32)
Since the discrete Fourier transform satisfies $\|\mathcal{F}y\|_{\infty}\leq\frac{1}{\sqrt{n}}\|y\|_{1}$ for $y\in\mathbb{C}^{n}$, we obtain
$\sqrt{p}\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}\leq\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}\,,$ (33)
which again splits the error into a time component and a frequency component.
It remains to obtain suitable decay of samples of $\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f}$ and $f-\mathcal{P}\!_{p}f$. We start with the frequency component $\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$. The observation $|\frac{k}{p}|\leq\frac{1}{2h}$, for $k\in[n]$, and assuming without loss of generality that $h\leq 1$ lead to
$\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$ $=\sup_{k\in[n]}\Big{|}\sum_{l\in\mathbb{Z}\setminus\{0\}}\hat{f}(\tfrac{k}{p}+h^{-1}l)\Big{|}\leq\sum_{|l|\geq\frac{1}{2h}}\sup_{\xi\in[0,1]}|\hat{f}(\xi+l)|\,.$ (34)
Since $\hat{f}\in W(\mathcal{C},\ell^{1})$, this sum tends to zero for $h\rightarrow 0$.
To estimate the time component $\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}$, we compute
$\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}=h\sum_{j\in[n]}\Big{|}\sum_{l\in\mathbb{Z}\setminus\{0\}}f(hj+pl)\Big{|}\leq h\sum_{j\in\mathbb{Z}\setminus[n]}|f(hj)|\,.$
Since $hj\in[l,l+1)$ occurs for at most $\lceil h^{-1}\rceil$ many integers $j$, we obtain
$\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}\leq h\lceil h^{-1}\rceil\sum_{|l|\geq\frac{p}{2}}\sup_{x\in[0,1]}|f(x+l)|\,.$
Since $f\in W(\mathcal{C},\ell^{1})$, the series tends to zero for $p\rightarrow\infty$. The factor $h\lceil h^{-1}\rceil\leq(1+h)$ does not cause any issues because $h$ is small, after all, we let $h\rightarrow 0$.
Since $\|y\|_{2}\leq\sqrt{n}\,\|y\|_{\infty}$ for $y\in\mathbb{C}^{n}$, Proposition 5.1 also implies the following $\ell^{2}$-error:
$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sqrt{h}\,E^{[n]}_{h}(f)\leq\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sqrt{h}\frac{\sqrt{n}}{\sqrt{p}}\,\sup_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}\rightarrow 0\,.$
for $f,\hat{f}\in W(\mathcal{C},\ell^{1})$.
In a similar spirit we can ask when the interpolation in (28) with the $B$-splines $B_{1}$ or $B_{2}$ converges to zero.
Proposition 5.3.
If $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then
$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\,\sup_{\xi\in\mathbb{R}}\Big{|}\hat{f}(\xi)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\Big{|}=0\,,\qquad i=1,2.$ (35)
A similar result under a stronger condition on $f$ and $\hat{f}$ and the choice $p=1/h=\sqrt{n}$ was derived in [16] for the piecewise linear interpolation with $B_{2}$. Proposition 5.3 improves the main result in [16] by extending the approximation result to a larger class of functions. For the rather non-trivial comparison of the condition $f,\hat{f}\in W(\mathcal{C},\ell^{1})$ to the one in [16] we refer to [17, Theorem 2]).
Proof 5.4.
As in the proof of Theorem 4.1 we split the approximation error into discretization, truncation, and approximation via the discrete Fourier transform as follows:
$\hat{f}(\xi)\,\overset{(a1)}{\approx}\,\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$ $\,\overset{(a2)}{\approx}\,\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$
$\,\overset{(a3)}{\approx}\,h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\,.$
To estimate the error in $(a1)$, we use the fact that the integer shifts of $B_{i}$, $i=1,2$, form a partition of unity, so that $\operatorname*{supp}(B_{i})\subseteq[-1,1]$ leads to
$\Big{|}\hat{f}(\xi)-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big{|}$ $=\Big{|}\sum_{k\in\mathbb{Z}}(\hat{f}(\xi)-\hat{f}(\tfrac{k}{p}))B_{i}(p\xi-k)\Big{|}$
$\leq\Big{|}\sum_{\begin{subarray}{c}k\in\mathbb{Z}\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}|\hat{f}(\xi)-\hat{f}(\tfrac{k}{p})|B_{i}(p\xi-k)\Big{|}$
$\leq\sup_{k\in\mathbb{Z}}\sup_{\xi\in\mathbb{R}:|\xi-\frac{k}{p}|\leq\frac{1}{p}}\Big{(}\sum_{k\in\mathbb{Z}}B_{i}(p\xi-k)\Big{)}$
$\leq\sup_{\xi,\eta\in\mathbb{R}:|\xi-\eta|<\frac{1}{p}}|\hat{f}(\xi)-f(\eta)|$
Since $f\in W(\mathcal{C},\ell^{1})$ is uniformly continuous and $p\rightarrow\infty$, we conclude that
$\lim_{p\rightarrow\infty}\sup_{\xi\in\mathbb{R}}\Big{|}\hat{f}(\xi)-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big{|}=0.$
For the error in $(a2)$, we use $\|B_{i}\|_{\infty}=1$ and find for all $\xi\in\mathbb{R}$ that
$\Big{|}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$ $-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big{|}=\Big{|}\sum_{k\in\mathbb{Z}\setminus[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big{|}$
$\leq\|B_{i}\|_{\infty}\sum_{\begin{subarray}{c}k\in\mathbb{Z}\setminus[n]\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}\big{|}\hat{f}(\tfrac{k}{p})\big{|}\leq\sum_{k\in\mathbb{Z}\setminus[n]}\big{|}\hat{f}(\tfrac{k}{p})\big{|}\,.$
Since $p\geq 1$ and $\frac{|k|}{p}\geq\frac{n}{2}\,\frac{1}{nh}=\frac{1}{2h}$ for $k\in[n]$, the last sum is estimated as in (34) by
$\sum_{k\in\mathbb{Z}\setminus[n]}\big{|}\hat{f}(\tfrac{k}{p})\big{|}\leq\sum_{|l|\geq\frac{1}{2h}}\sup_{\xi\in[0,1]}|\hat{f}(\xi+l)|\to 0\,.$
Since $\hat{f}\in W(\mathcal{C},\ell^{1})$, we obtain
$\lim_{h\rightarrow 0}\sup_{\xi\in\mathbb{R}}\Big{|}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big{|}=0\,.$
In order to estimate $(a3)$, we compute
$\bigg{|}\sum_{k\in[n]}\Big{(}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{)}B_{i}(p\xi-k)\bigg{|}$
$\qquad\leq\sup_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}\,\Big{|}\sum_{k\in[n]}B_{i}(p\xi-k)\Big{|}$
$\qquad\leq\sup_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}\,,$
were we have used again that the integer translates of the cardinal $B$-splines form a partition of unity. According to Proposition 5.1 this term vanishes when $p\rightarrow\infty$ and $h\rightarrow 0$.
The final error is the sum of the three errors (a1), (a2), and (a3).
Appendix A Time frequency symmetry of the error functional
We prove the time-frequency symmetry of the error $E^{[n]}_{h}$.
Lemma A.1.
If $p=nh$, then
$E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}}),\qquad f,\hat{f}\in W(\mathcal{C},\ell^{1})\,.$ (36)
Hence, the time frequency decomposition of the error matches the time-frequency symmetry of $E^{[n]}_{h}$.
Proof A.2.
To verify (36), we recall that $(\bar{\hat{f}})\,\widehat{}=\bar{f}$ and that $\mathcal{F}$ is unitary, so that
$E^{[n]}_{p^{-1}}(\bar{\hat{f}})$ $=\|(\overline{\widehat{f_{h,n}}})\,\widehat{}-\mathcal{F}\bar{\hat{f}}_{\frac{1}{p},n}\|=\|\bar{f}_{h,n}-\mathcal{F}\bar{\hat{f}}_{\frac{1}{p},n}\|$
=*$=\|\mathcal{F}^{*}\bar{f}_{h,n}-\bar{\hat{f}}_{\frac{1}{p},n}\|=E^{[n]}_{p^{-1}}(\bar{\hat{f}})\,.$
The discrete Fourier transform also satisfies $\overline{\mathcal{F}y}=\mathcal{F}^{*}\bar{y}$, where conjugation is meant entry-wise. Conjugation does not affect the norm, so that we obtain (36).

### v2

@@
 We derive exact error estimates for this procedure in terms of the decay and smoothness of $f$.
-The analysis provides a new recipe of how to relate the number of samples, the sampling interval, and the grid size.
+The analysis provides an asymptotically optimal recipe of how to relate the number of samples, the sampling interval, and the grid size.
 keywords:
@@
 The fast Fourier transform (FFT) is widely used in the applied sciences for the numerical approximation of the Fourier transform
-$\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}x\xi}\mathrm{d}x$ (1)
+(1) $\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}x\xi}\mathrm{d}x$
 from sampled values of $f$.
@@
 As the FFT is a fast algorithm that computes the discrete Fourier transform, we are actually asking: how well does the discrete Fourier transform approximate the Fourier transform [11]?
+1.1 Questions
 Engineers and numerical practitioners compute the Fourier transform of a function $f$ on the real line as follows:
   • (i)
-    Sample $f$ on an interval of length $p$ at $n$ equispaced points $hj$, for $j\in\mathbb{Z}$, $-\frac{n}{2}<j\leq\frac{n}{2}$, with step size $h$.
+    Sample $f$ on an interval $(-\frac{p}{2},\frac{p}{2}]$ of length $p$ at $n$ equispaced points $hj$, for $j\in\mathbb{Z}$, $\frac{n}{2}<j\leq\frac{n}{2}$, with step size $h$ and $p=hn$.
   • (ii)
-    It is then taken for granted that the scaled discrete Fourier transform of length $n$ provides an approximation for the values of the continuous Fourier transform of $f$ at $k/p$, i.e.,
-    $\hat{f}\big{(}\tfrac{k}{p}\big{)}\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\,,\qquad k\in\mathbb{Z},\quad-\frac{n}{2}<k\leq\frac{n}{2}\,.$ (2)
-This practice is usually unquestioned and works like a charm. It is motivated by following approximations with quadrature and truncation,
-$\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}\xi x}\mathrm{d}x\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\,.$ (3)
-For $\xi=\frac{k}{p}=\frac{k}{hn}$, the right-hand side is the scaled discrete Fourier transform (2).
-Yet one must ask what is the precise relation between the computed values $h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}$ and the samples $\hat{f}(\frac{k}{p})$ of the actual Fourier transform.
+    Approximate the Fourier transform of $f$ at $\tfrac{k}{p}$ by
+    (2) $\hat{f}\big(\tfrac{k}{p}\big)\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\,,\qquad k\in\mathbb{Z},\quad-\frac{n}{2}<k\leq\frac{n}{2}\,.$
+The right-hand side is the discrete Fourier transform of the sampled vector $\big(f(hj)\big)_{j}$ of size $n$.
+By choosing $n$ to be a power of $2$ or by zero-padding, the discrete Fourier transform can be computed with the FFT algorithm in $\mathcal{O}(n\log n)$ operations.
+Since the Fourier transform is a fundamental computational step in a huge number of applications in signal processing, acoustics, medical imaging, the numerical solution of partial differential equations, in quantum mechanics, etc., the availability of a fast algorithm is central in scientific computing.
+Therefore much effort has gone into a speed up and efficient implementations of this algorithm [25, Chapter 12].
+The FFT is hailed as “the most important numerical algorithm of our life-time” [29], is included in the ten most important numerical algorithms of the 20th century [9], and remains important in the age of data science and artificial intelligence.
+Grounded on the overwhelming success of the FFT in approximating the Fourier transform, the approximation (2) is usually taken for granted.
+The FFT is used as a black box, it works superbly, and it always seems to work.
+Qualitatively, the approximation (2) is easy to understand. It is motivated by discretizing the integral and then truncating the infinite series as follows:
+(3) $\hat{f}(\xi)=\int_{-\infty}^{\infty}f(x)e^{-2\pi\mathrm{i}\xi x}\mathrm{d}x\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\approx h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\xi hj}\,.$
+For $\xi=\frac{k}{hn}=\frac{k}{p}$ one then obtains (2).
+Although this approximation is immensely successful, one must ask what is the precise relation between the computed values $h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}$ and the samples $\hat{f}(\frac{k}{p})$ of the actual Fourier transform.
 Ideally one can provide reasonable error estimates for this approximation, and we formulate two questions:
   • (Q1)
-    The numerical procedure depends on the number of samples $n$, the step size $h$, and the length $p$ of the sampled interval via the identity $p=hn$.
-How should $h,n$ and $p$ be chosen in dependence of the function class of $f$?
+    The numerical procedure depends on the number of samples $n$, the step size $h$, and the length $p$ of the sampled interval where $p=hn$.
+How should $h,n$, and $p$ be chosen dependent on the function class of $f$?
   • (Q2)
-    What can be said about the asymptotic decay of the deviation in (2) when $n,p\rightarrow\infty$ and $h\to 0$?
-State-of-the-art. To the best of our knowledge, there are surprisingly few investigations and even fewer quantitative error estimates in the literature.
-The “Owner’s Manual for the Discrete Fourier Transform” [4, Section 6] contains a substantial chapter on the error of the discrete Fourier transform.
-There the error is quantified for rather special function classes, e.g., functions with compact support or functions on $\mathbb{R}$ with exponential decay with some type of finite order of smoothness conditions [4, Theorem 6.6] that are rarely met, see also our Section 3.5.
-[…]
-Since the right-hand side of (2) approximates the integral $\int_{-p/2}^{p/2}f(x)e^{-2\pi ikx/n}\mathrm{d}x$ by means of a Riemann sum or the trapezoidal rule, one should mention the recent error analyses [7, 11, 23, 25, 24] for the trapezoidal rule as relevant in this context.
-However, these error estimates require exponential decay and analyticity and do not even mention the Fourier transform.
-In this paper we derive estimates for the error between the Fourier transform on the real line and the discrete Fourier transform for a large class of functions with polynomial or sub-exponential decay and finite smoothness.
-Our results confirm that, even under such very mild assumptions on the decay and smoothness, the folklore procedure works well and is successful.
-Our analysis also provides the new insight that the optimal spacing should depend on the decay of $f$ and $\hat{f}$.
-In view of (Q1), we identify an optimal relation between the step size $h$, the number of samples $n$, and the length of the sampling interval $p$ (the step size $\frac{1}{p}$ in frequency).
-To answer (Q2), we derive precise error estimates with explicit constants for function classes that are significantly larger than exponentially decaying, analytic functions.
-These error estimates are confirmed by numerical simulations utilizing the FFT which yield the precise asymptotics predicted by our theoretical estimates.
-Our contribution. We proceed with a formal presentation of the numerical procedure and an explicit exposition of our results.
+    What can be said about error rates and the asymptotic decay of the deviation in (2) when $n\rightarrow\infty$, $p\rightarrow\infty$, and $h\to 0$?
+We will answer these questions for several classes of functions that are described by (i) their decay property and (ii) by their smoothness or by the decay of their Fourier transform.
+The decay of $f$ determines the truncation error in (3), and the decay of $\hat{f}$ determines the discretization error in (3).
+1.2 State-of-the-art
+The case of functions with compact support is related to the case of approximating Fourier coefficients of a periodic function and is fully understood.
+Roughly, if $f$ has compact support and is $m$-times differentiable, then the pointwise error scales like $n^{-m}$ in terms of length of the FFT, or like $h^{m}$ in terms of the grid size.
+This is made precise by Epstein in [11], whose title we borrowed, for the approximation of Fourier coefficients of periodic functions, and by Briggs and Henson in their “Owner’s Manual for the Discrete Fourier Transform” [5, Section 6] for the Fourier transform of compactly supported functions on $\mathbb{R}$.
+For $\operatorname*{supp}f\subseteq[-\frac{p}{2},\frac{p}{2}]$ and $h=\frac{p}{n}$, these results are optimal and come with explicit constants.
+The article [2] computes an exact formula for the relative pointwise error for the Fourier transform of a compactly supported $B$-spline of order $k$ (called “canonicalk-k functions”) and shows that in the limit $n\to\infty$ the relative error is the same for $k+1$-times differentiable functions with compact support, however, no error estimates are given.
+To the best of our knowledge, there are surprisingly few investigations and even fewer quantitative error estimates in the literature that go beyond compactly supported functions.
+For functions with unbounded support the important cases of exponential decay and analyticity of $f$ are covered by the work of Briggs and Henson [5, Chapter 6] and of Stenger [28, Chapter 3.3].
+The case of general input in [5, Thm.
+6.6] assumes that $f$ satisfies $\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}<\infty$ and is $m$-times differentiable satisfying some stringent boundary conditions at $\pm p/2$.
+Then for some constants $c_{1},c_{2}$ the pointwise error is
+(4) $\Big|\hat{f}\big(\tfrac{k}{p}\big)-h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi i\frac{kj}{n}}\Big|\leq c_{1}e^{-rp/2}e^{-rp/n}+c_{2}h^{m},\qquad-\frac{n}{2}<k\leq\frac{n}{2}\,.$
+The decay of order $h^{m}$ is as expected, and the implicit assumption of bounded variation [5] even leads to $h^{m+1}$.
+The problem, however, is in the boundary conditions.
+In the generic case, when the values of $f$ at the endpoints of the sampled interval differ, i.e., $f(-p/2)\neq f(p/2)$, then the predicted error is only of the order $h$.
+For instance, for the exponential function $f(x)=e^{-2\pi|x|}$ [5] obtains the correct error estimate with $h^{2}$, but for the shifted function $e^{-2\pi|x-1|}$ the predicted error is only $h^{1}$, which does not match the numerical observations.
+This phenomenon seems to be a limitation of the proof technique in [5].
+It is one of our contributions to obtain the correct error estimate without any boundary conditions.
+For more discussion see our Section 4.7.
+In the context of numerical sinc methods Stenger [28, Chapter 3.3] uses the cardinal series to approximate an exponentially decaying function $f$ whose Fourier transform also decays exponentially (to be precise, he uses a slightly stronger condition formulated with a Hardy space) and shows that the pointwise error satisfies
+$\Big|\hat{f}\big(\tfrac{k}{p}\big)-h\sum_{-\frac{n}{2}<j\leq\frac{n}{2}}f(hj)e^{-2\pi i\frac{kj}{n}}\Big|\leq Ce^{-sn^{1/2}}\,$,
+with suitable constants $C,s>0$. In addition, [28, Chapter 3.3] contains error estimates for the $L^{p}$-error for interpolated versions of the Fourier transform.
+Since the right-hand side of (2) approximates the integral $\int_{-p/2}^{p/2}f(x)e^{-2\pi\mathrm{i}xk/p}\mathrm{d}x$ by means of a Riemann sum, one should mention the recent error analyses [10, 14, 30, 32, 31] for the trapezoidal rule as relevant in this context although they do not mention the Fourier transform.
+These error estimates assume exponential decay and analyticity of $f$.
+By contrast, the recent results for the trapezoidal rule in [21, Prop.
+4.2, Thm 4.5] do not require analyticity, but use Gaussian decay and finite Sobolev smoothness of order $m$.
+For $n$ samples of $f$ and the choice $p=2\sqrt{\frac{2}{1-\epsilon}m\log n}$ the error is of the order $n^{-m}(\log n)^{m/2+1/4}$.
+In [8] the trapezoidal rule is used for error estimates of approximations of the Fourier transform of analytic functions via doubly exponential transforms.
+These are rather different from the usual approximation (2), but yield almost exponential convergence.
+An interesting variation was studied by Auslander and Grünbaum [1].
+They consider functions $f\in L^{2}(\mathbb{R})$ and approximate averaged values of $\hat{f}$ from local averages of $f$ via the discrete Fourier transform.
+For Gaussian smoothing they derive an explicit formula for the sharp constant for the resulting error estimate and then plot the results of numerical simulations for FFTs up to $n=100$ samples.
+Asymptotics for large $n$ is not treated.
+2 Our contributions
+In this paper we derive estimates for the error between the Fourier transform on the real line and the discrete Fourier transform for a significantly broader class of functions than previously considered in the literature.
+Specifically, we study functions that decay only polynomially in time and in frequency.
+Our results confirm in a quantitative fashion that, even under mild assumptions on the decay and smoothness (decay in frequency), the standard approximation procedure with the FFT works well and is successful.
+Our analysis also provides the new insight how the optimal spacing should depend on the decay of $f$ and $\hat{f}$.
+As an answer to Question (Q1), we identify an optimal relation between the step size $h$, the number of samples $n$, and the length $p$ of the sampling interval.
+To answer Question (Q2), we derive precise error estimates with explicit constants for function classes that are significantly larger than exponentially decaying or analytic functions.
+The optimality of the error estimates is confirmed by numerical simulations with the FFT that yield the precise asymptotics predicted by theoretical results.
+We believe that these error estimates are best possible, and since all constants are explicit, these estimates cover not only the asymptotic regime, but also hold for small $n$.
+We proceed with the detailed exposition of our results.
+2.1 Error measures
 We first introduce the appropriate notation. For $n\in\mathbb{N}$, set
-$[n]:=\{j\in\mathbb{Z}:-\tfrac{n}{2}<j\leq\tfrac{n}{2}\}\,.$
+$n]:=\{j\in\mathbb{Z}:-\tfrac{n}{2}<j\leq\tfrac{n}{2}\}\,.$
 The discrete Fourier transform $\operatorname{\mathcal{F}}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ of $y=(y_{j})_{j\in[n]}\in\mathbb{C}^{n}$ is
-$\operatorname{\mathcal{F}}\!y=\bigg{(}\frac{1}{\sqrt{n}}\sum_{j\in[n]}y_{j}e^{-2\pi\mathrm{i}\frac{kj}{n}}\bigg{)}_{k\in[n]}\,.$ (4)
+(5) $\operatorname{\mathcal{F}}\!y=\bigg(\frac{1}{\sqrt{n}}\sum_{j\in[n]}y_{j}e^{-2\pi\mathrm{i}\frac{kj}{n}}\bigg)_{k\in[n]}\,.$
 Define the scaled sampling of a function $f$ with step size $h$ and length $n$ as
 $f_{h,n}:=\Big(\sqrt{h}f(hj)\Big)_{j\in[n]}\in\mathbb{C}^{n}\,.$
-Throughout the text we use the relation $p=hn$ between the interval length $p$, the step size $h$ and the number of samples $n$.
-Our aim is to bound the approximation error
-$E^{[n]}_{h}(f):=\biggl{(}\frac{1}{p}\sum_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}^{2}\bigg{)}^{1/2}=\big{\|}\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\big{\|}\,.$ (5)
-The normalization of $f_{h,n}$ and the factor $\frac{1}{p}$ in the definition of the error may seem artificial, but they are justified and natural in view of the time-frequency symmetry of the error
-$E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}})\,.$
-See Appendix A for the proof. The symmetry between $f$ and $\hat{f}$ will also appear in the error estimates.
-The following error estimate is a simple consequence of our main result for polynomial decay.
+Throughout the text we use the relation $p=hn$ between the interval length $p$, the step size $h$, and the number of samples $n$.
+We will deal mainly with the $\ell^{2}$-error averaged over the sampled interval. This approximation error is defined as
+(6) $E^{[n]}_{h}(f):=\biggl(\frac{1}{p}\sum_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|^{2}\bigg)^{1/2}=\big\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\big\|\,.$
+The use of the $\ell^{2}$-error is natural in our context, since the discrete Fourier transform is unitary on $\mathbb{C}^{n}$ with respect to the $\ell^{2}$-norm.
+The normalization of $f_{h,n}$ and the factor $\frac{1}{p}$ are motivated by the remarkable time-frequency symmetry of the error
+[…]
+see Lemma 4.1.
+This symmetry aligns well with the mathematical framework of time-frequency analysis in [17] in which both time and frequency are treated as equally important dimensions for understanding signals.
+The symmetry between time and frequency also manifests in our error estimates, and we believe that the unified time-frequency perspective is a key reason why our estimates are sharp.
+2.2 Polynomial decay and finite smoothness
+Our main theorem will be proved for general decay conditions.
+We first discuss the special case of functions of polynomial decay and finite smoothness.
+As remarked in  [5], finite smoothness is often replaced by polynomial decay of the Fourier transform, which we will do here.
+The first error estimate is a simple consequence of our main result (Theorems 4.7 and 4.6) and is the most relevant special case.
 Theorem 2.1.
 Let $a,b>1$.
-If $f$ and $\hat{f}$ are continuous and satisfy polynomial decay of the form [^1] We write $\lesssim$ if the left-hand-side is bounded by a constant times the right-hand-side.
+If $f$ and $\hat{f}$ are continuous and satisfy polynomial decay of the form $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}<\infty$, then, for all $\alpha<a-\frac{1}{2}$ and $\beta<b-\frac{1}{2}$, there is a constant $c>0$ that is independent of $h$, $n$, $p$ such that
+(7) $E^{[n]}_{h}(f)\leq c\,(p^{-\alpha}+h^{\beta})\,.$
+This statement can be read on several levels.
+  1. (i)
+    It confirms the users’ experience that the discrete Fourier transform indeed approximates the continuous Fourier transform.
+  2. (ii)
+    Theorem 2.1 offers guarantees of how well the FFT approximates the Fourier transform by quantifying the approximation error in terms of the decay of $f$ and $\hat{f}$.
+  3. (iii)
+    The theorem provides a hint on how to improve the accuracy of the numerical computations. This is a new insight with possibly practical consequences.
+    If no a priori knowledge of $f$ is given, then the $n$ samples of $f$ are usually chosen on an interval of length $p=\sqrt{n}$ with grid size $h=1/{\sqrt{n}}$.
+In this case the error decays like $E^{[n]}_{h}(f)=\mathcal{O}(n^{-\min(\alpha,\beta)/2})$.
+However, with some knowledge about these decay rates, one may balance the two terms in (7) so that both terms contribute equally***We write $\lesssim$ if the left-hand-side is bounded by a constant times the right-hand-side.
 If $\lesssim$ and $\gtrsim$ both hold, then we write $\asymp$.
-The dependency of the constants on other parameters shall be clarified or is clear from the context.
-$|f(x)|\lesssim(1+|x|)^{-a}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$, then for all $\alpha<a-\frac{1}{2}$, $\beta<b-\frac{1}{2}$
-$E^{[n]}_{h}(f)\lesssim p^{-\alpha}\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,,$
-where the constants in $\lesssim$ are independent of $h$, $n$, $p$.
-The error bound is balanced if $p^{-\alpha}\asymp h^{\beta}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$, and for this choice $E^{[n]}_{h}(f)\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}$.
-This is the optimal choice for the relation between $h$, $p$, and $n$, see Corollary 3.9.
-Our numerical simulations in Section 2 using the FFT suggest that the bounds in Theorem 1.1 are sharp.
-For the approximation procedure to make sense, we need the Fourier transform to be defined pointwise and continuous, so that $f\in L^{1}(\mathbb{R})$ is a reasonable minimal assumption.
+The dependency of the constants on other parameters shall be clarified or is clear from the context., i.e., $p^{-\alpha}\asymp h^{\beta}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$.
+For this choice the error is
+    (8) $E^{[n]}_{h}(f)\leq cn^{-\frac{\alpha\beta}{\alpha+\beta}}$
+    This observation will become relevant when the decay rates $\alpha,\beta$ are known and the difference $|\alpha-\beta|$ is large.
+The choice $p^{-\alpha}\asymp h^{\beta}$ then leads to a significantly more accurate numerical approximation of the Fourier transform using the same number of samples.
+Conversely, a given target level of accuracy can be achieved with significantly fewer samples.
+See Section 3.
+    We also note that this choice leads to the asymptotically optimal relation between $h$, $p$, and $n$, see Corollary 4.11.
+  4. (iv)
+    In Corollary 4.12 we will determine the constants involved in the balancing $p^{-\alpha}\asymp h^{\beta}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$, and we will determine the constant $c$ in (8) explicitly in terms of $\alpha,\beta$ and a norm on $f$.
+We emphasize that the ultimate version of the error estimate is a non-asymptotic result that may yield relevant information about the error even for small $n$.
+Our numerical simulations in Section 3 confirm that the bounds in Theorem 2.1 are sharp up to some infinitesimal margin.
+Smoothness instead of Fourier decay
 One may argue that we also make certain assumptions on $\hat{f}$, which is the very object we want to compute.
-If one does not like the polynomial decay condition of $\hat{f}$ in Theorem 1.1, one can replace it by a stronger condition on smoothness of $f$.
-Precisely, if the derivatives $f^{(l)}$ are in $L^{1}(\mathbb{R})$ for $l=0,\dots,b$ with $b\in\mathbb{N}$, then it follows that $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$.
-Sub-exponential decay in time and in frequency occurs importantly in the theory of test functions and distributions (under the name of Gelfand-Shilov spaces or ultra test functions) [3].
-In that case, we offer the following version.
+The polynomial decay condition of $\hat{f}$ in Theorem 2.1 can be replaced by a stronger condition on smoothness of $f$.
+Precisely, if the derivatives $f^{(l)}$ are in $L^{1}(\mathbb{R})$ for $l=0,\dots,b$ with $b\in\mathbb{N}$, then it follows that $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}<\infty$.
+Using this observation, we obtain a slightly weaker, but equally useful version of Theorem 2.1.
+Corollary 2.2.
+Let $a,m>1$.
+Assume that $|f(x)|\lesssim(1+|x|)^{-a}$ and that $f$ has $m$ derivatives in $L^{1}(\mathbb{R})$.
+Then for $0<\alpha<a-1/2$ and $\beta<m-1/2$ there is a constant $c>0$ independent of $h$, $n$, $p$ such that
+(9) $E^{[n]}_{h}(f)\leq c\,(p^{-\alpha}+h^{\beta})\,.$
+2.3 Sub-exponential decay
+Exponential and sub-exponential decay in time and in frequency occur importantly in the theory of test functions and distributions (under the name of Gelfand-Shilov spaces or ultra test functions) [3].
+The strongest form is exponential decay for which error estimates have been treated before.
+We offer the following version of the error.
 Theorem 2.3 (sub-exponential decay).
 Let $r,s>0$ and $0<\alpha,\beta\leq 1$.
-If $f$ and $\hat{f}$ are continuous and satisfy the decay conditions $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim e^{-s|x|^{\beta}}$, then for all $r^{\prime}<r$,s<$s^{\prime}<s$,
-$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}+e^{-s^{\prime}(2h)^{-\beta}}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|e^{s|\xi|^{\beta}}\,.$
+If $f$ and $\hat{f}$ are continuous and satisfy the decay conditions $\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|e^{s|\xi|^{\beta}}<\infty$, then for all $r^{\prime}<r$ and $s^{\prime}<s$, there is a constant $c>0$ that is independent of $h$, $n$, $p$ such that
+$E^{[n]}_{h}(f)\leq c\,(e^{-r^{\prime}(\frac{p}{2})^{\alpha}}+e^{-s^{\prime}(2h)^{-\beta}})\,.$
 For $r^{\prime}\to r$ and $s^{\prime}\to s$ the error is balanced if $e^{-r(\frac{p}{2})^{\alpha}}\asymp e^{-s(2h)^{-\beta}}$, which is satisfied for $h=n^{-\frac{\alpha}{\alpha+\beta}}\left(\tfrac{s}{r}\right)^{\frac{1}{\alpha+\beta}}2^{\frac{\alpha-\beta}{\alpha+\beta}}$.
 This condition on $h$ is the optimal choice for the relation between $h$, $p,$ and $n$, see Corollary 4.18.
-We also provide bounds for mixed decay.
+For the parameters $\alpha=\beta=1$, the assumption says that both $f$ and $\hat{f}$ decay exponentially. The choice $p=\sqrt{n}$ and $h=1/{\sqrt{n}}$ then yields the error estimate
+$E^{[n]}_{h}(f)\leq c\,e^{-d\sqrt{n}}\,$,
+with a constant $d$ depending on $r,s,\alpha,\beta$. Up to possibly different constants, this root-exponential convergence is compatible with Stenger’s result [28].
+2.4 Mixed conditions
+We finally provide bounds for mixed conditions when $f$ possesses exponential decay, but finite smoothness.
 Theorem 2.4 (mixed decay).
 Let $r>0$ and $0<\alpha\leq 1$.
-If $f$ and $\hat{f}$ are continuous and satisfy $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$,
-$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,.$ (6)
-For all practical purposes, the length of the sampled interval $p$ is sufficiently large, so that $e^{-r^{\prime}(\frac{p}{2})^{\alpha}}$ is negligible (and may for instance even be close to machine precision).
+If $f$ and $\hat{f}$ are continuous and satisfy $\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}<\infty$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$, there is a constant $c>0$ that is independent of $h$, $n$, $p$ such that
+(10) $E^{[n]}_{h}(f)\leq c\,(e^{-r^{\prime}p/2}+h^{\beta})\,.$
+For all practical purposes, the length of the sampled interval $p$ is sufficiently large, so that $e^{-r^{\prime}p/2}$ is negligible (and usually close to machine precision).
 Then the error is dominated by the polynomial term $h^{\beta}$.
-In the main theoretical part in Section 3, we will derive more general results by quantifying the decay of a function by means of its membership in a Wiener amalgam space.
-This is the canonical function class when sampling is involved [9, 10, 12].
+The above estimate implies a pointwise error of $h^{\beta}$ for $\beta<b-1/2$.
+The error estimate (4) from [5] induces a bound for the $\ell^{2}$-error $E^{[n]}_{h}(f)$ on the order of $h^{\beta}$ for $\beta=b-\frac{1}{2}$, however, under additional, stringent boundary conditions.
+For comparison, we formulate the error estimate for the known case of functions with compact support.
+Corollary 2.5.
+Assume that $f$ has compact support in $\frac{p}{2},\frac{p}{2}]$ with $p=hn$ and is $m$-times differentiable, then for every $\epsilon>0$ there is a constant $c=c(\epsilon,p)$, such that
+$E^{[n]}_{h}(f)\leq c\,n^{-(m-1/2-\epsilon)}\,.$
+We note that the pointwise estimate $n^{-m}$ from [11] translates into the estimate $E^{[n]}_{h}(f)\lesssim n^{-(m-1/2)}$.
+Thus except for the additional $\epsilon$ in the exponent, we recover the well-known results for compactly supported functions.
+The presence of $\epsilon$ is plausible, because the assumption of differentiability is slightly stronger than the polynomial decay of $\hat{f}$, and we derive Corollary 2.5 from that weaker assumption (and thus obtain a weaker conclusion).
+2.5 Interpolation
 Software to plot the discrete Fourier transform commonly performs interpolation of the computed data vector $\operatorname{\mathcal{F}}\!f_{h,n}$.
+The implicit question is then what the samples of $\hat{f}(k/p)$ or their approximations say about the global Fourier transform $\hat{f}$.
 In order to simulate this process, we use three representative interpolation procedures and derive bounds on their deviation from $\hat{f}$ on the real line.
-One of them is based on the cardinal sine function $\operatorname*{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$ that is widely used in digital signal processing.
+One of them is based on the cardinal sine function $\operatorname*{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$ that is widely used in time-frequency analysis, digital signal processing, and numerical analysis [6, 24, 27, 28].
 The computation of $\operatorname{\mathcal{F}}\!f_{h,n}$ gives rise to the approximation of $\hat{f}$ on the real line by
 $\hat{f}(\xi)\approx\sqrt{p}\sum_{k\in[n]}(\mathcal{F}f_{h,n})(k)\operatorname*{sinc}(p\xi-k)=h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\xi-k)\,.$
-So the error is then $\|\hat{f}-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\cdot-k)\|_{L^{2}(\mathbb{R})}$.
-We will show in Section 5, Theorem 4.1, that this error obeys the same bounds as in Theorems 1.1, 1.2, and 1.3.
-The proof uses the estimates for $E^{[n]}_{h}(f)$.
-Outline: Section 2 is dedicated to numerical experiments using the FFT and fully supports the theoretical results of Theorems 1.1 and 1.3.
+We will consider the $L^{2}$ norm, this is, we estimate
+$\|\hat{f}-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\cdot-k)\|_{L^{2}(\mathbb{R})}\,$,
+and will show in Section 5, Theorem 5.1, that this error obeys exactly the same bounds as $E_{h}^{[n]}(f)$ in Theorems 2.1, 2.3, and 2.4.
+2.6 Minimal requirements for convergence
+Whereas in most applications and in scientific computing it is important to understand how well the FFT approximates the Fourier transform and how fast the convergence happens, it is a fundamental question whether and when the approximations converge at all.
+Mathematically, we search for minimal conditions on a function.
+To the best of our knowledge, this problem has not been addressed in the literature.
+For the initial formulation, we use the pointwise error and weak polynomial decay conditions.
+Proposition 2.6.
+If $f$ and $\hat{f}$ satisfy the mild decay conditions $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{1+\epsilon}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{1+\epsilon}<\infty$ for some $\epsilon>0$, then
+(11) $\lim_{p\to\infty}\lim_{h\to 0}\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|=0\,.$
+Clearly, these decay conditions are the weakest possible. If $\epsilon=0$, then $f$ need not be integrable and its Fourier transform may not even be defined pointwise.
+In Section 6 we will formulate a more general, but technical condition that is related to the validity of the Poisson summation formula, and we will also treat the convergence of the interpolation.
+2.7 Other error measures and noisy samples
+We will be exclusively interested in the $\ell^{2}$-approximation error $E^{[n]}_{h}(f)$, that is the error made by replacing the Fourier transform by the FFT.
+Other error norms or the influence of other effects, such as rounding or distortions, can be derived from the main estimates.
+(i) Pointwise error estimates follow easily because
+(12) $\max_{k\in[n]}\Big|\hat{f}\Big(\tfrac{k}{p}\Big)-h\sum_{j\in[n]}f(hj)e^{-2\pi i\frac{kj}{n}}\Big|\leq\sqrt{p}E^{[n]}_{h}(f)\,$,
+and other $\ell^{q}$-norms can be handled similarly. As an example, under the decay assumptions of Theorem 2.1, (8) implies the pointwise error estimate
+$\max_{\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\lesssim n^{-\frac{\beta}{\alpha+\beta}(\alpha-\frac{1}{2})}\,$,
+where the last inequality follows for the choice $p^{-\alpha}\asymp n^{-\frac{\alpha\beta}{\alpha+\beta}}$.
+We do not claim that the bounds for the pointwise error are optimal.
+Our tools are designed for $E^{[n]}_{h}(f)$ and rely heavily on its time-frequency symmetry, which is not shared by the pointwise error.
+(ii) Noise, contaminated function evaluation, or round-off errors are important sources of errors in scientific computing.
+In each case, instead of the correct samples $f(jh)$ the samples of a distorted function $\widetilde{f}$ or noisy samples $\widetilde{f}(hj)=f(hj)+\epsilon_{j}$ are used as input of the FFT.
+However, since the approximation (2) is linear and the FFT is unitary, the effect of these errors is benign and remains what it was.
+Formally, using $\widetilde{f}$ and its samples $\widetilde{f}_{h,n}$ instead of $f_{h,n}$ as input of the FFT, the resulting error is
+(13) $\big\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!\widetilde{f}_{h,n}\big\|\leq\big\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\big\|+\big\|\operatorname{\mathcal{F}}\!f_{h,n}-\operatorname{\mathcal{F}}\!\widetilde{f}_{h,n}\big\|=E^{[n]}_{h}(f)+\big\|(\widetilde{f}-f)_{h,n}\big\|\,.$
+Thus the error splits into the approximation error $E^{[n]}_{h}(f)$, which we are concerned with, and the norm of the initial noise.
+Consequently, all error bounds remain valid for the distorted function values after adding the noise term $\big\|(\widetilde{f}-f)_{h,n}\big\|$ to the original bound.
+2.8 Methods
+As in [5] we will use the Poisson summation formula to decompose the approximation error into a time and a frequency component.
+This allows us to estimate each component separately.
+In order to make this approach precise, we need to sample a function on a grid and to apply a pointwise version of Poisson’s summation formula.
+The novelty of our approach is (i) the systematic exploitation of the time-frequency symmetry of the problem, and (ii) the use of so-called Wiener amalgam spaces to describe sampling and decay properties.
+This class is the canonical function class when sampling is involved [12, 13, 15] and is widely used in Fourier analysis.
+We believe that it is also very useful and convenient in numerical analysis.
+Outline
+Section 3 is dedicated to numerical simulations to illustrate that the theoretical results of Theorems 2.1 and 2.4 are best possible.
 The full versions of the theoretical results on the approximation rates for spaces of polynomial and sub-exponential decay in time and in frequency are derived in Section 4.
-The approximation of $\hat{f}$ on the real line is discussed in Section 4.
-We investigate minimal requirements on $f$ and $\hat{f}$ for convergence in Section 5.
-2 Numerical experiments for polynomial decay
-Since our priority is the practical impact of the error estimates, we first offer numerical evidence that the error bounds of Theorem 1.1 are sharp for polynomial weights.
+The approximation of $\hat{f}$ on the real line via interpolation is discussed in Section 5.
+In Section 6 we investigate under which minimal conditions the error tends to zero.
+Acknowledgements
+The authors thank Norbert Kaiblinger for valuable suggestions that helped improving the manuscript.
+We would like to express our gratitude to the referees for their meticulous reading and detailed comments on the manuscript.
+Their suggestions have helped us immensely to improve the manuscript.
+A.
+K.
+is supported by the FWF project DISCO (PAT4780023).
+3 Numerical simulations
+We back up the theoretical results with numerical evidence that the polynomial error bounds in Theorem 2.1 are sharp, up to some infinitesimal margin.
+It is important to emphasize that the following numerical examples are not intended to represent typical functions for a specific application.
+The sole purpose of this section is to illustrate that the assumptions and error rates associated with polynomial decay cannot, in general, be improved.
+Strictly speaking, our theoretical results allow for arbitrary $\alpha<a-\frac{1}{2}$, while the numerics suggest that the rate $\alpha=a-\frac{1}{2}$ still holds.
+This is what we mean by ‘sharp up to some infinitesimal margin’.
+This numerical verification of optimality is nontrivial: it requires functions $f$ for which both $f$ and its Fourier transform $\hat{f}$ can be evaluated with high accuracy, and for which the decay rates in both time and frequency are precisely known.
+For functions with a standard closed-form expression, at least one of these properties typically fails.
+We therefore construct suitable functions through a more refined approach below.
 3.1 Polynomial decay in time and in frequency
 In view of Theorem 2.1, we construct a family $f^{a,b}$ of functions with exact polynomial decay $a$ in time and $b$ in frequency such that each function and its Fourier transform can be numerically evaluated with sufficient accuracy.
-We use linear combinations of shifts of the cardinal B-splines defined by $B_{1}:={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$, and
+We use infinite linear combinations of shifts of the cardinal B-splines defined by $B_{1}:={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$, and
 $B_{b+1}:=B_{b}*B_{1},\qquad b=1,2,\ldots\,.$
-Each $B_{b}$ is a piecewise polynomial function of degree $b-1$ that is $b-2$-times continuously differentiable with support $[-\frac{b}{2},\frac{b}{2}]$, see [20] and also [19, Section 9.1].
+Each $B_{b}$ is a piecewise polynomial function of degree b1b-1 that is b2b-2-times continuously differentiable with support $\frac{b}{2},\frac{b}{2}]$, and its Fourier transform is $\widehat{B_{b}}(\xi)=\big(\tfrac{\sin\pi\xi}{\pi\xi}\big)^{b}=\operatorname*{sinc}^{b}(\xi)$, see [26] and also [24, Section 9.1].
 Consider $(c_{k})_{k\in\mathbb{Z}}\subseteq\mathbb{C}$ given by its nonzero entries
 $c_{0}=\frac{1}{2},\qquad c_{k}=\frac{1}{k\pi\mathrm{i}},\text{ for }k\in 2\mathbb{Z}+1\,$,
-so that $c_{2k}=0$. We study the family of functions
-$f^{a,b}(x):=\sum_{k\in\mathbb{Z}}c^{a}_{k}\,B_{b}(x-k),\qquad a,b=1,2,\ldots\,,$ (7)
-which consists of locally finite sums. Therefore the point evaluations can be evaluated accurately in numerical experiments, and they satisfy $|f^{a,b}(x)|\lesssim(1+|x|)^{-a}$.
-To compute the Fourier transform of $f^{a,b}$, we define the Fourier series $u_{a}(\xi):=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}$.
-As $\widehat{B_{b}}=\operatorname*{sinc}^{b}$, we observe
+so that $c_{2k}=0$. For integer parameters $a,b\in\mathbb{N}$ we study the family of functions
+(14) $f^{a,b}(x):=\sum_{k\in\mathbb{Z}}c^{a}_{k}\,B_{b}(x-k),\qquad a,b=1,2,\ldots\,.$
+Every $f^{a,b}$ is a locally finite sum, therefore its point evaluations can be computed accurately in numerical experiments, and they satisfy $\sup_{x\in\mathbb{R}}|f^{a,b}(x)|(1+|x|)^{a}<\infty$.
+To compute the Fourier transform of $f^{a,b}$, we define the Fourier series $u_{a}(\xi):=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}$ and obtain
 $\widehat{f^{a,b}}(\xi)$ $=\operatorname*{sinc}(\xi)^{b}\,\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}=\operatorname*{sinc}(\xi)^{b}u_{a}(\xi)\,.$
-The estimate $|\operatorname*{sinc}^{b}(\xi)|\lesssim(1+|\xi|)^{-b}$ implies $|\widehat{f^{a,b}}(\xi)|\lesssim(1+|\xi|)^{-b}$, thus $f^{a,b}$ satisfies the assumptions of Theorem 1.1.
+The estimate $|\operatorname*{sinc}^{b}(\xi)|\lesssim(1+|\xi|)^{-b}$ implies $\sup_{\xi\in\mathbb{R}}|\widehat{f^{a,b}}(\xi)|(1+|\xi|)^{b}<\infty$, thus $f^{a,b}$ satisfies the assumptions of Theorem 2.1.
 By a short computation we find $u_{1}(\xi)={\bf 1}_{(-\frac{1}{2},0)}(\xi)+\frac{1}{2}{\bf 1}_{\{-\frac{1}{2},0\}}(\xi)$ for $\xi\in[-\frac{1}{2},\frac{1}{2})$, and then $u_{a+1}$ is the cyclic convolution of $u_{a}$ and $u_{1}$, i.e.,
 $u_{a+1}(\xi)=\int_{\mathbb{R}/\mathbb{Z}}u_{a}(\eta)u_{1}(\xi-\eta)\mathrm{d}\eta,\qquad a=1,2,\ldots\,.$
-For $a=2,3$ we have $u_{2}(\xi)=|\xi|$ $u_{a}$ for $\xi\in[-\frac{1}{2},\frac{1}{2})$, and $u_{3}(\xi)=\operatorname*{sign}(\xi)\xi^{2}-\frac{1}{2}\xi+\frac{1}{8}$.
-Thus, we can evaluate the point evaluations of $f^{a,b}$ and $\widehat{f^{a,b}}$ exactly, both analytically and numerically, and the decay is labeled by $a$ and $b$.
+For $a=2,3,4,5$, tedious calculations lead to periodic functions with period $1$ whose values on [1/2,1/2)[-1/2,1/2) are given by
+$u_{2}(\xi)$ $=|\xi|,$
+$u_{3}(\xi)$ $=-\operatorname*{sign}(\xi)\xi^{2}+\frac{1}{2}\xi+\frac{1}{8},$
+$u_{4}(\xi)$ $=2|\xi|^{3}/3-\xi^{2}/2+1/12,$
+$u_{5}(\xi)$ $=-\operatorname*{sign}(\xi)\xi^{4}/3+\xi^{3}/3-\xi/24+1/32\,.$
+Thus, we can compute the point evaluations of $f^{a,b}$ and $\widehat{f^{a,b}}$ exactly, both analytically and numerically, and the decay parameters are precisely $a$ and $b$.
 For the numerical simulations we compute the samples $\widehat{f^{a,b}}(k/p)$ and compare them with the discrete Fourier transform of a sampled version of $f^{a,b}$.
-Since the error estimates of Theorem 1.1 work for every parameter $\alpha<a-1/2$ and $\beta<b-1/2$, we will plot the error curves with the values $\alpha=a-\frac{1}{2}$ and $\beta=b-\frac{1}{2}$ and compare our numerical observations with the error rate $n^{-\frac{\alpha\beta}{\alpha+\beta}}$ for $h=n^{-\frac{\alpha}{\alpha+\beta}}$.
-The infinitesimal difference between $\alpha<a-1/2$ and $\alpha=a-1/2$ is not visible.
-To compute the discrete Fourier transform, we choose $n=2^{l}$, $l=10,\ldots,20$ in Theorem 1.1 and then apply the FFT to the samples $f^{a,b}_{h,n}$.
-The FFT is taken from the Julia package FFTW.jl with $64$-bit floating point arithmetics.
-The package provides a binding to the C library FFTW [13].
-The approximation rates, predicted by Theorem 1.1, are $n^{-\frac{\alpha\beta}{\alpha+\beta}}\asymp h^{\beta}$, see Table 1 for some combinations of $a$ and $b$.
-To stick to $L^{1}(\mathbb{R})$ functions, we skip $a=1$.
+To compute the discrete Fourier transform, we choose $n=2^{l}$, $l=10,\ldots,20$ and then apply the FFT to the samples $f^{a,b}_{h,n}$.
+The FFT is taken from the Julia package FFTW.jl that provides a binding to the C library FFTW [16].
+Since the error estimates of Theorem 2.1 work for every parameter $\alpha<a-1/2$ and $\beta<b-1/2$, we choose $\alpha=a-\frac{1}{2}$ and $\beta=b-\frac{1}{2}$.
+We compare the results of the numerical experiments with the error rates $n^{-\frac{\alpha\beta}{\alpha+\beta}}$ predicted by Theorem 2.1 for $h=n^{-\frac{\alpha}{\alpha+\beta}}$ that are listed in Table 1.
 $a$ $b$ $\alpha=a-\frac{1}{2}$ $\beta=b-\frac{1}{2}$ $h=n^{-\frac{\alpha}{\alpha+\beta}}$ $p=nh=n^{\frac{\beta}{\alpha+\beta}}$ $E^{[n]}_{h}(f^{a,b})\asymp h^{\beta}=n^{-\frac{\alpha\beta}{\alpha+\beta}}$
-$2$ $2$ $3/2$ $3/2$ $n^{-1/2}$ $n^{1/2}$ $n^{-3/4}$
-$2$ $3$ $3/2$ $5/2$ $n^{-3/8}$ $n^{5/8}$ $n^{-15/16}$
-$2$ $4$ $3/2$ $7/2$ $n^{-3/10}$ $n^{7/10}$ $n^{-21/20}$
-$3$ $2$ $5/2$ $3/2$ $n^{-5/8}$ $n^{3/8}$ $n^{-15/16}$
-$3$ $3$ $5/2$ $5/2$ $n^{-1/2}$ $n^{1/2}$ $n^{-5/4}$
-$3$ $4$ $5/2$ $7/2$ $n^{-5/12}$ $n^{7/12}$ $n^{-35/24}$
+22 22 $3/2$ $3/2$ $n^{-1/2}$ $n^{1/2}$ $n^{-3/4}$
+22 33 $3/2$ $5/2$ $n^{-3/8}$ $n^{5/8}$ $n^{-15/16}$
+22 44 $3/2$ $7/2$ $n^{-3/10}$ $n^{7/10}$ $n^{-21/20}$
+33 22 $5/2$ $3/2$ $n^{-5/8}$ $n^{3/8}$ $n^{-15/16}$
+33 33 $5/2$ $5/2$ $n^{-1/2}$ $n^{1/2}$ $n^{-5/4}$
+33 44 $5/2$ $7/2$ $n^{-5/12}$ $n^{7/12}$ $n^{-35/24}$
 Table 1: List of the relevant values of Theorem 2.1 for $f^{a,b}$, so that $E^{[n]}_{h}(f^{a,b})$ is bounded by $h^{\beta}=p^{-\alpha}=n^{-\frac{\alpha\beta}{\alpha+\beta}}$.
-Our numerical experiments as illustrated in Figure 1 align with the theoretical bounds in Theorem 1.1 and suggest that they are sharp.
+Our numerical experiments as illustrated in Figure 1 align perfectly with the theoretical bounds in Theorem 2.1 as presented in Table 1 and suggest that they are sharp up to some infinitesimal margin.
 Figure 1: Logarithmic plots of $E^{[n]}_{h}(f^{a,b})$ for various $a$ and $b$.
 Points are computed numerically for $n=2^{l}$, $l=10,\ldots,20$.
-Reference lines represent the theoretical slopes from Theorem 1.1 listed in Table 1.
+Reference lines represent the theoretical slopes predicted by Theorem 2.1 and listed in Table 1.
 Remark 3.1.
-The closed form expressions for $u_{a}(\xi)=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi\mathrm{i}k\xi}$ are key to our numerical results: Without a closed form expression of $u_{a}$ the series in (7) would need to be truncated for evaluation.
-The truncation error is accompanied by competing floating point errors due to the summation.
-We found it difficult to identify a sufficiently accurate truncation that yields reliable numerical estimates for $E^{[n]}_{h}(f^{a,b})$.
-2.2 Polynomial decay in frequency
-To test Theorem 1.3, we consider $f(x)=e^{-2\pi|x|}$.
-Its Fourier transform is $\hat{f}(\xi)=\frac{1}{\pi(1+\xi^{2})}$, so that $r=2\pi$, $\alpha=1$, and $|\widehat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b=2$.
-For increasing values of $n$ in numerical experiments with the FFT, we choose successively $p=n^{1/2}$, $p=6n^{1/10}$, and $p=10$, so that $h=n^{-1/2}$, $h=6n^{-9/10}$, and $h=10n^{-1}$, respectively.
-Note that $p$ is sufficiently large, so that the error bound in Theorem 1.3 is expected to be governed by the polynomial term for a large range of $n$.
-Hence, we expect to observe the decay rate $h^{\frac{3}{2}}$, which is $n^{-\frac{3}{4}}$, $n^{-\frac{27}{10}}$, and $n^{-\frac{3}{2}}$, respectively.
-Figure 2 shows the results of our numerical experiments.
-Again, they match the theoretical prediction perfectly.
+Let us revisit the construction of $f^{a,b}$ to highlight why it is so remarkable that we are able to perform numerical experiments.
+The B-splines $B_{b}$ and their Fourier transforms $\operatorname*{sinc}^{b}$ can be evaluated accurately.
+Because B-splines have compact support, the infinite series defining $f^{a,b}$ reduces to a finite sum when evaluated at a finite number of points.
+Moreover, the careful choice of coefficients in this expansion ensures that the Fourier transform of $f^{a,b}$ admits a closed-form expression and enables the accurate numerical evaluation.
+This is a rare situation, since most functions do not satisfy these properties, and meaningful numerical experiments depend critically on them.
+In particular, the closed-form expressions for $u_{a}(\xi)=\sum_{k\in\mathbb{Z}}c^{a}_{k}e^{-2\pi ik\xi}$ are essential to our numerical results.
+Without them, the series in (14) would need to be truncated, introducing truncation errors.
+Summing many terms also risks floating-point inaccuracies, reducing the reliability of computed values of $E^{[n]}_{h}(f^{a,b})$.
+The availability of a closed-form expression for $u_{a}$ eliminates these sources of numerical error.
+Choice of $p$ and $h$
+To highlight the importance of a good choice of $p$ and $h$, we consider the function $f^{2,4}$, $f^{2,5}$, $f^{3,5}$, $f^{5,2}$, whose decay in time differs from the decay in frequency.
+We compare two parameter selections: the optimal choice suggested by Theorem 2.1, versus the standard choice $p=\sqrt{n}$ and $h=1/\sqrt{n}$ [4, 20].
+As shown in Figures 2, 3, 4, and 5, the optimal parameters yield an error decay rate consistent with our theoretical predictions and clearly outperform the standard choice.
+From the plots one can draw valuable information about the number of samples required to achieve a target accuracy.
+For instance, by drawing a horizontal line in Figure 2 at the error $E_{h}^{[n]}(f^{2,4})=2^{-15}\approx 3\cdot 10^{-5}$ we find that $n=2^{12}$ samples are required for the optimal choice of $h,n,p$, whereas $2^{17}$ samples are required for the choice $p=h^{-1}=\sqrt{n}$.
+Even for an accuracy of only five digits, this is an enormous saving!
+Figure 2: For $f^{2,4}$, we compare the error curve for the standard choice $p=\sqrt{n}$, $h=1/\sqrt{n}$ with the error curve for the improved choice $p=n^{7/10}$, $h=n^{-3/10}$.
+The reference lines indicate the theoretical error decay rates: $n^{-3/4}$ for the standard choice and $n^{-21/20}$ for the optimal one.
+Figure 3: Comparison of error curves for $f^{2,5}$.
+Figure 4: Comparison of error curves for $f^{3,5}$.
+Figure 5: Comparison of error curves for $f^{5,2}$.
+3.2 Exponential decay in time and polynomial decay in frequency
+To demonstrate that Theorem 2.4 is best possible, we consider the exponential functions
+$f_{\delta}(x)=e^{-2\pi|x-\delta|},$
+for a shift parameter $\delta\in\mathbb{R}$. Its Fourier transform is
+$\hat{f_{\delta}}(\xi)=\frac{e^{-2\pi i\delta\xi}}{\pi(1+\xi^{2})}.$
+Thus, $f_{\delta}$ decays exponentially in space and polynomially in frequency.
+This corresponds to the parameters in Theorem 2.4 being $r=2\pi$, $\alpha=1$, and $\sup_{\xi\in\mathbb{R}}|\hat{f_{\delta}}(\xi)|(1+|\xi|)^{b}<\infty$ with $b=2$.
+When $p$ is sufficiently large, the error in Theorem 2.4 is dominated by the polynomial decay over a wide range of $n$. For numerical experiments using the FFT with $\delta=0$, we consider the choices
+$p=n^{1/2},\quad p=6n^{1/10},\quad\text{and}\quad p=10\,$,
+corresponding to $h=n^{-1/2}$, $h=6n^{-9/10}$, and $h=10n^{-1}$, respectively. The expected convergence rates are $h^{3/2}$, which translate into
+$n^{-3/4},\quad n^{-27/20},\quad\text{and}\quad n^{-3/2}\,.$
+Figure 6 confirms these predictions with excellent agreement between theory and experiment for all shift parameters $\delta$ (the plots look identical for $0\leq\delta\leq 1$).
 Figure 6: Logarithmic plots of $E^{[n]}_{h}(f)$ for $f(x)=e^{-2\pi|x|}$.
@@
 We choose $p=n^{\frac{1}{2}}$, $p=6n^{1/10}$, and $p=10$.
-Reference lines represent the theoretical slope $h^{\beta}$ with $\beta=\frac{3}{2}$ that leads to $n^{-\frac{3}{4}}$, $n^{-\frac{27}{20}}$, and $n^{-\frac{3}{2}}$, respectively.
+Reference lines represent the theoretical slope $h^{\beta}$ with $\beta=\frac{3}{2}$ that lead to $n^{-\frac{3}{4}}$, $n^{-\frac{27}{20}}$, and $n^{-\frac{3}{2}}$, respectively.
 Points align well with theoretical predictions.
+This example is also considered in a case study in [5].
+For $\delta=0$ the pointwise error is calculated to be of the order $n^{-2}$.
+However, as soon as $\delta\neq 0$ and $p=10$, say, the required condition $f_{\delta}(p/2)=f_{\delta}(-p/2)$ in [5] is violated, and their bounds yield only $n^{-1}$ for the maximum norm, and hence only $n^{-1/2}$ for the $2$-norm as used in $E^{[n]}_{h}(f)$, whereas both Theorem 2.4 and our numerical experiments consistently yield the error rate $n^{-3/2}$.
+This further supports the sharpness of our theoretical bounds.
 4 Approximation rates
-In this section we derive general and explicit bounds on the approximation error $E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|$ in (5) by decomposing the error into a time and a frequency component that are estimated separately.
-These components arise from approximating $\hat{f}(\tfrac{k}{p})$ first by a Riemannian sum and then truncating it as in (3),
-$\hat{f}\big{(}\tfrac{k}{p}\big{)}\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}\approx h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=\sqrt{hn}(\operatorname{\mathcal{F}}\!f_{h,n})(k)\,.$ (8)
-The left-hand side is the sampled (continuous) Fourier transform, the right-hand side is its approximation by the discrete Fourier transform.
-The idea is to estimate the discretization error $\hat{f}\big{(}\tfrac{k}{p}\big{)}-h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$ by the decay of $\hat{f}$, and the truncation $h\sum_{j\in\mathbb{Z}\setminus[n]}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$ by the decay of $f$.
+In this section we prove Theorems 2.1, 2.3, and 2.4 from the introduction.
+All these statements follow from a single theorem by feeding it with various decay conditions.
+Specifically, we derive error bounds for the averaged $\ell^{2}$-error $E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|$.
+This error exhibits a striking symmetry between time and frequency, which will be reflected in the symmetry of the error estimates.
+Lemma 4.1.
+If $p=nh$ and $f,\hat{f}\in L^{1}(\mathbb{R})$, then
+[…]
+Proof 4.2.
+To verify (15), we recall that $(\bar{\hat{f}})\,\widehat{}=\bar{f}$ and that $\mathcal{F}$ is unitary, so that
+[…]
+The discrete Fourier transform also satisfies $\mathcal{F}^{*}\bar{y}=\overline{\mathcal{F}y}$, where complex conjugation is meant entry-wise.
+Therefore, we derive $E^{[n]}_{p^{-1}}(\bar{\hat{f}})=\|\overline{\mathcal{F}f_{h,n}}-\bar{\hat{f}}_{\frac{1}{p},n}\|$.
+Complex conjugation does not affect the norm, so that the latter coincides with $\|\mathcal{F}f_{h,n}-\hat{f}_{\frac{1}{p},n}\|=E^{[n]}_{h}(f)$.
+Thus, we expect good bounds for $E^{[n]}_{h}(f)$ to reflect the same symmetry between $(f,h)$ and $(\hat{f},p^{-1})$.
 4.1 Error decomposition into time and frequency components
-The continuous functions on the real line are denoted by $\mathcal{C}=\mathcal{C}(\mathbb{R})$.
-We consider the Wiener amalgam space $W(\mathcal{C},\ell^{1})$ that consists of all continuous functions $f$ such that
-$\|f\|_{W(\mathcal{C},\ell^{1})}:=\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|$ (9)
-is finite, cf.
+We derive general and explicit bounds on the approximation error $E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|$ in (6) by decomposing the error into a time and a frequency component that are estimated separately.
+These components arise from approximating $\hat{f}(\tfrac{k}{p})$ first by a Riemann sum and then truncating it as in (3),
+(16) $\hat{f}\big(\tfrac{k}{p}\big)\approx h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}\approx h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=\sqrt{hn}(\operatorname{\mathcal{F}}\!f_{h,n})(k)\,.$
+The left-hand side is the sampled (continuous) Fourier transform, the right-hand side is its approximation by the discrete Fourier transform. The idea is to estimate the discretization error
+$\hat{f}\big(\tfrac{k}{p}\big)-h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$
+by the decay of $\hat{f}$, and the truncation error
+$h\sum_{j\in\mathbb{Z}\setminus[n]}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$
+by the decay of $f$.
+4.2 Wiener amalgam spaces
+The treatment of the error makes use of sampling a function on a grid and a strong pointwise version of the Poisson summation formula.
+To handle these, we introduce the Wiener amalgam spaces that are tailored to these objectives.
+The continuous functions on the real line are denoted by $\mathcal{C}=\mathcal{C}(\mathbb{R})$. The space $W(\mathcal{C},\ell^{1})$ consists of all continuous functions $f$ such that
+(17) $\|f\|_{W(\mathcal{C},\ell^{1})}:=\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|$
+is finite.
+Wiener amalgam spaces were introduced by N.
+Wiener [33] and have become convenient and important function spaces in Fourier analysis, cf.
 [12, 13, 15].
-If $f\in W(\mathcal{C},\ell^{1})$, then the sampling operator $f\mapsto(f(hj))_{j\in\mathbb{Z}}\in\ell^{1}(\mathbb{Z})$ is bounded for all $h>0$, and consequently the $p$-periodization
+The norm in (17) is easy to understand.
+A function $f$ belongs to $W(\mathcal{C},\ell^{1})$, if it is continuous and is majorized by an integrable step function with jumps at the integers.
+See Figure 7.
+Figure 7: Visualization of the construction of the Wiener amalgam norm of a continuous function $f$.
+In the context of the discretization of the Fourier transform this space arises naturally, because both sampling and periodization are well defined.
+If $f\in W(\mathcal{C},\ell^{1})$, then the sampling operator $f\mapsto(f(hj))_{j\in\mathbb{Z}}$ maps $W(\mathcal{C},\ell^{1})$ to $\ell^{1}(\mathbb{Z})$ for all $h>0$, and Here $\lceil\cdot\rceil$ denotes the ceiling function.
+$\sum_{j\in\mathbb{Z}}|f(jh)|\leq\lceil\tfrac{1}{h}\rceil\|f\|_{W(\mathcal{C},\ell^{1})}\,$,
+Consequently the $p$-periodization
 $\mathcal{P}\!_{p}f(x)=\sum_{l\in\mathbb{Z}}f(x+pl)$
 converges absolutely and uniformly for every $p>0$.
-For $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, the Poisson summation formula $\sum_{l\in\mathbb{Z}}f(x+pl)=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi i\frac{k}{p}x}$ holds pointwise for all $x\in\mathbb{R}$, cf.
-[14, Lemma 4].
+Furthermore, if $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then the Poisson summation formula
+$\sum_{l\in\mathbb{Z}}f(x+pl)=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi i\frac{k}{p}x}$
+holds pointwise for all $x\in\mathbb{R}$, cf. [18, Lemma 4]. No simple, larger space seems to be known on which the Poisson summation formula holds pointwise for all periods $p$.
 The following identity provides the foundation of our approach and will facilitate the error decomposition in a time and a frequency component.
@@
 If $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then, for $k\in[n]$,
-$h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}=h\sum_{j\in[n]}(\mathcal{P}\!_{p}f)(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=(\mathcal{P}\!_{h^{-1}}\hat{f})(\tfrac{k}{p})\,.$ (10)
-We may also write the second identity in terms of $(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}=\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}$.
-Proof 3.2.
-The first identity is the Poisson summation formula applied to the $h^{-1}$-periodization of $\hat{f}$.
-For the second identity, we partition $\mathbb{Z}$ into residue classes modulo $n$ and use $p=hn$.
-Then
+(18) $h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}=h\sum_{j\in[n]}(\mathcal{P}\!_{p}f)(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}=(\mathcal{P}\!_{h^{-1}}\hat{f})(\tfrac{k}{p})\,.$
+We may also write the second identity as $\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}=(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}$.
+Proof 4.4.
+For the first identity, we partition $\mathbb{Z}$ into residue classes modulo $n$ and use $p=hn$. Then
 $h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}\frac{k}{p}hj}$ $=h\sum_{j\in[n]}\sum_{l\in\mathbb{Z}}f(hj+hnl)e^{-2\pi\mathrm{i}\frac{k}{hn}(hj+hnl)}$
@@
 $=h\sum_{j\in[n]}\mathcal{P}\!_{p}f(hj)e^{-2\pi\mathrm{i}\frac{k}{n}j}\,.$
+The second identity in (18) follows from the Poisson summation formula applied to the $h^{-1}$-periodization of $\hat{f}$,
+$(\mathcal{P}\!_{h^{-1}}\hat{f})(\tfrac{k}{p})$ $=h\sum_{j\in\mathbb{Z}}\hat{\hat{f}}(hj)e^{2\pi\mathrm{i}hj\frac{k}{p}}$
+$=h\sum_{j\in\mathbb{Z}}f(hj)e^{-2\pi\mathrm{i}hj\frac{k}{p}}\,.$
+This is again the left-hand-side of (18).
 Lemma 4.3 enables a reinterpretation of the two step approximation strategy via quadrature and truncation (16).
@@
 Symbolically,
-$\hat{f}_{\frac{1}{p},n}\approx(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}=\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}\approx\operatorname{\mathcal{F}}\!f_{h,n}\,.$ (11)
-More precisely, we split the pointwise error in two components,
-$|\frac{1}{\sqrt{p}}\hat{f}(\tfrac{k}{p})-\operatorname{\mathcal{F}}\!f_{h,n}(k)|\leq\frac{1}{\sqrt{p}}|\hat{f}(\tfrac{k}{p})-\mathcal{P}\!_{h^{-1}}\hat{f}(\tfrac{k}{p})|+|\operatorname{\mathcal{F}}\Big{(}(\mathcal{P}\!_{p}f-f)_{h,n}\Big{)}(k)|\,.$ (12)
-Taking norms and using the unitarity of $\operatorname{\mathcal{F}}$ we arrive at
-$\begin{split}E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|\leq&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|\operatorname{\mathcal{F}}\Big{(}(\mathcal{P}\!_{p}f-f)_{h,n}\Big{)}\|\\ =&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(\mathcal{P}\!_{p}f-f)_{h,n}\|\,.\end{split}$ (13)
-On the right hand side of this inequality we have (i) eliminated $\operatorname{\mathcal{F}}$ and (ii) split the error in a time component and a frequency component.
-Note that the symmetry of this decomposition matches the symmetry of the error $E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}})$, cf.
-Appendix A.
-To derive approximation rates, we must ensure that the samples of $f-\mathcal{P}\!_{p}f$ and $\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f}$ decay sufficiently fast.
-This can be accomplished by specifying decay conditions on $f$ and $\hat{f}$.
-3.2 General weights
-We quantify the decay of a function by the use of a weight function $v:\mathbb{R}\rightarrow(0,\infty)$. We consider weights $v$ that satisfy
-$v\text{ is even: }v(-x)=v(x)\,,$ (14a)
-$v\text{ is nondecreasing, i.e., }v(x)\leq v(y),\text{ for }|x|\leq|y|\,,$ (14b)
-$v\text{ is submultiplicative, i.e., }v(x+y)\leq v(x)v(y)\text{ for }x,y\in\mathbb{R}\,,$ (14c)
-$v\text{ is inverse square-summable, }\sum_{l\in\mathbb{Z}}{|v(l)|^{-2}}<\infty\,.$ (14d)
-The standard examples of weights that satisfy conditions (14a) to (14d) are the polynomial weights $v_{\alpha}=(1+|x|)^{\alpha}$ for $\alpha>\frac{1}{2}$, and the (sub-)exponential weights $v_{r,\alpha}=e^{r|x|^{\alpha}}$, for $0<\alpha\leq 1$ and $r>0$.
+(19) $\hat{f}_{\frac{1}{p},n}\approx(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}=\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}\approx\operatorname{\mathcal{F}}\!f_{h,n}\,.$
+We therefore split the pointwise error in two components,
+(20) $|\frac{1}{\sqrt{p}}\hat{f}(\tfrac{k}{p})-\operatorname{\mathcal{F}}\!f_{h,n}(k)|\leq\frac{1}{\sqrt{p}}|\hat{f}(\tfrac{k}{p})-\mathcal{P}\!_{h^{-1}}\hat{f}(\tfrac{k}{p})|+|\operatorname{\mathcal{F}}\big((\mathcal{P}\!_{p}f-f)_{h,n}\big)(k)|\,.$
+[…]
+(21) $\begin{split}E^{[n]}_{h}(f)=\|\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}\|\leq&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|\operatorname{\mathcal{F}}(\mathcal{P}\!_{p}f-f)_{h,n}\|\\ =&\,\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(\mathcal{P}\!_{p}f-f)_{h,n}\|\,.\end{split}$
+[…]
+To handle the sampling of a function on a grid simultaneously with its decay properties, we introduce another family of Wiener amalgam spaces.
+We first quantify the decay of a function by the use of a weight function $v:\mathbb{R}\rightarrow(0,\infty)$.
+We consider weights $v$ that satisfyThe condition (22d) is crucial, whereas the conditions (22a), (22b), and (22c) can be weakened at the expense of more complicated constants in the error estimates.
+(22a) v $(x),\displaystyle v\text{ is even: }v$(-x)=v(x)\,,
+[…]
+(22c) $v\text{ is submultiplicative, i.e., }v(x+y)\leq v(x)v(y)\text{ for }x,y\in\mathbb{R}\,$,
+(22d) $v^{-1}\text{ is square-summable, }\sum_{l\in\mathbb{Z}}{|v(l)|^{-2}}<\infty\,.$
+The standard examples of weights that satisfy conditions (22a) to (22d) are the polynomial weights $v_{\alpha}(x)=(1+|x|)^{\alpha}$ for $\alpha>\frac{1}{2}$, and the (sub-)exponential weights $v_{r,\alpha}(x)=e^{r|x|^{\alpha}}$, for $0<\alpha\leq 1$ and $r>0$.
 Associated to $v$ are the weighted sequence spaces
@@
 and the weighted Lebesgue space $L^{2}_{v}:=L^{2}_{v}(\mathbb{R})$ of all functions $f$ such that $\|f\|_{L^{2}_{v}}=\|fv\|_{L^{2}}$ is finite.
-For a constant weight $v\equiv 1$ we omit the subscript and write $\ell^{p}$ and $L^{p}$.
+For the constant weight $v\equiv 1$ we omit the subscript and write $\ell^{p}$ and $L^{p}$.
 The Wiener amalgam space $W(\mathcal{C},\ell^{q}_{v})$ consists of all continuous functions $f$ such that
-$\|f\|_{W(\mathcal{C},\ell^{q}_{v})}:=\bigg{(}\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{q}|v(l)|^{q}\bigg{)}^{1/q}$ (15)
-is finite, cf. [9, 10, 12].
-If $v$ satisfies (14d) then $\ell^{2}_{v}\subseteq\ell^{1}$, and we derive the continuous embeddings $W(\mathcal{C},\ell^{2}_{v})\subseteq W(\mathcal{C},\ell^{1})\subseteq L^{1}(\mathbb{R})$.
-In order to bound the right hand side of the error functional $E^{[n]}_{h}(f)$ in (13), we need to estimate the deviation of $f$ and $\hat{f}$ from their respective periodizations.
+(23) $\|f\|_{W(\mathcal{C},\ell^{q}_{v})}:=\bigg(\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{q}|v(l)|^{q}\bigg)^{1/q}$
+is finite. For the theory and applications of Wiener amalgam spaces we recommend  [12, 13, 15].
+If $v$ satisfies (22d) then $\ell^{2}_{v}\subseteq\ell^{1}$, and we obtain the continuous embeddings $W(\mathcal{C},\ell^{2}_{v})\subseteq W(\mathcal{C},\ell^{1})\subseteq L^{1}(\mathbb{R})$.
+4.3 Towards the main error estimates
+In order to bound the right-hand side of the error functional $E^{[n]}_{h}(f)$ in (21), we need to estimate the deviation of $f$ and $\hat{f}$ from their respective periodizations.
 We derive such bounds in terms of the function
-$\Phi_{v}(p):=\bigg{(}2\sum_{m=0}^{\infty}|v(pm+\tfrac{p}{2})|^{-2}\bigg{)}^{1/2}\,.$ (16)
-Observe that $\lim_{p\to\infty}\Phi_{v}(p)=0$ always holds for a weight $v$ satisfying (3.2).
-For polynomial and sub-exponential weights we obtain quantitative bounds on the decay of $\Phi_{v}(p)$ in Lemma 3.7 and in Lemma 3.13.
+(24) $\Phi_{v}(p):=\bigg(2\sum_{m=0}^{\infty}|v(pm+\tfrac{p}{2})|^{-2}\bigg)^{1/2}\,.$
+Observe that $\lim_{p\to\infty}\Phi_{v}(p)=0$ always holds for a weight $v$ satisfying (22).
+For polynomial and sub-exponential weights we will obtain quantitative bounds on the decay of $\Phi_{v}(p)$ in Lemma 4.9 and in Lemma 4.16.
 Lemma 4.5.
-If $f\in W(\mathcal{C},\ell^{2}_{v})$, and $v$ satisfies (3.2) then
-$\|f-\mathcal{P}\!_{p}f\|_{L^{2}(-\frac{p}{2},\frac{p}{2})}$ $\leq v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})},$ (i)
-$\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ $\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$ (ii)
+If $f\in W(\mathcal{C},\ell^{2}_{v})$ for a weight $v$ satisfying (22), then
+(i) $\|f-\mathcal{P}\!_{p}f\|_{L^{2}(-\frac{p}{2},\frac{p}{2})}$ $\leq v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})},$
+(ii) $\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ $\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$
 Part (i) is used only in Section 5 on interpolation, but Part (ii) is crucial for bounds on $E^{[n]}_{h}$.
-Proof 3.4.
-(i) We bound $(f-\mathcal{P}\!_{p}f)(x)$ by the Cauchy-Schwartz inequality,
-$\bigg{|}\sum_{0\neq l\in\mathbb{Z}}f(x+pl)\bigg{|}^{2}\leq\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}\sum_{0\neq l\in\mathbb{Z}}\left|f(x+pl)\right|^{2}|v(x+pl)|^{2}\,.$ (17)
+Proof 4.6.
+(i) We bound $(f-\mathcal{P}\!_{p}f)(x)=\sum_{0\neq l\in\mathbb{Z}}f(x+pl)$ by the Cauchy-Schwarz inequality,
+(25) $\bigg|\sum_{0\neq l\in\mathbb{Z}}f(x+pl)\bigg|^{2}\leq\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}\sum_{0\neq l\in\mathbb{Z}}\left|f(x+pl)\right|^{2}|v(x+pl)|^{2}\,.$
 Since $v$ is even (22a) and nondecreasing (22b), for $|x|\leq\frac{p}{2}$, the first sum can be majorized by
-$\begin{split}\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}&\leq\sum_{0\neq m\in\mathbb{Z}}|v(p|m|-\tfrac{p}{2})|^{-2}\\ &\leq 2\sum_{m=1}^{\infty}|v(pm-\tfrac{p}{2})|^{-2}=|\Phi_{v}(p)|^{2}\,.\end{split}$ (18)
+(26) $\begin{split}\sum_{0\neq m\in\mathbb{Z}}|v(x+pm)|^{-2}&\leq\sum_{0\neq m\in\mathbb{Z}}|v(p|m|-\tfrac{p}{2})|^{-2}\\ &\leq 2\sum_{m=1}^{\infty}|v(pm-\tfrac{p}{2})|^{-2}=|\Phi_{v}(p)|^{2}\,.\end{split}$
 Integration of (25) yields
@@
 $\leq|\Phi_{v}(p)|^{2}\sum_{j\in\mathbb{Z}}\left|f(hj)\right|^{2}|v(hj)|^{2}\,.$
-[…]
+Since $v$ satisfies (22a) – (22c), we derive
 $\sum_{j\in\mathbb{Z}}\left|f(hj)\right|^{2}|v(hj)|^{2}$ $\leq\lceil h^{-1}\rceil\sum_{l\in\mathbb{Z}}\sup_{x\in[l,l+1]}\left|f(x)\right|^{2}|v(|l|+1)|^{2}$
@@
 The observation $h\lceil h^{-1}\rceil\leq(1+h)$ concludes the proof.
+4.4 The main theorem
 After these preparations we can now formulate our main result.
-This is an error estimate for the approximation of the Fourier transform by the discrete Fourier transform with very general conditions on the decay of the function and its Fourier transform.
+This is an error estimate for the approximation of the Fourier transform by the discrete Fourier transform under very general conditions on the decay of the function and its Fourier transform.
 Theorem 4.7.
-Let $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$. If the weights $v,w$ satisfy (3.2), then
-$E^{[n]}_{h}(f)\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+w(1)\Phi_{w}(h^{-1})(1+\tfrac{1}{p})^{\frac{1}{2}}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (19)
-Note that for $0<h\leq 1\leq p$ the first term does not depend on $h$ and the second term not on $p$. Therefore
-$E^{[n]}_{h}(f)\leq 2^{\frac{1}{2}}v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+2^{\frac{1}{2}}w(1)\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (20)
+Let $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$. If the weights $v,w$ satisfy (22), then
+(27) $E^{[n]}_{h}(f)\leq v(1)\Phi_{v}(p)(1+h)^{\frac{1}{2}}\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+w(1)\Phi_{w}(h^{-1})(1+\tfrac{1}{p})^{\frac{1}{2}}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
+For $0<h\leq 1\leq p$, we obtain
+(28) $E^{[n]}_{h}(f)\leq 2^{\frac{1}{2}}v(1)\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+2^{\frac{1}{2}}w(1)\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,$,
+so that the first term does not depend on $h$ and the second term not on $p$.
 We now prove Theorem 4.7.
-Proof 3.6.
+Proof 4.8.
+Recall the bound $E^{[n]}_{h}(f)\leq\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ in (21).
 We apply Lemma 4.5 (ii) to both $f$ and $\hat{f}$.
-Recall the bound $E^{[n]}_{h}(f)\leq\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|+\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ in (13).
 Lemma 4.5 estimates $\|(f-\mathcal{P}\!_{p}f)_{h,n}\|$ and yields the first term of the error in (27).
-Applying Lemma 3.3 to $\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|$ by interchanging the roles of $f$, $p$, and $h$ with $\hat{f}$, $h^{-1}$, and $p^{-1}$, yields the second term.
+Applying Lemma 4.5 to $\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|$ and interchanging the roles of $f$, $p$, and $h$ with $\hat{f}$, $h^{-1}$, and $p^{-1}$, yields the second term.
 This concludes the proof.
 Theorem 4.7 asserts a bound on the error $E^{[n]}_{h}(f)$ in terms of $\Phi_{v}(p)$ and $\Phi_{w}(h^{-1})$.
-For polynomial and sub-exponential weights the error estimates can be made much more explicit.
+For useful conclusions we need to unravel the error function $\Phi_{v}$, and we will do this for polynomial and sub-exponential weights.
+The resulting error estimates can then be made completely explicit.
 4.5 Polynomial weights
-In this section we consider the polynomial weights $v_{\alpha}(x)=(1+|x|)^{\alpha}$.
-To bound $\Phi_{v_{\alpha}}(p)$ in (16), we make use of the Hurwitz zeta function $\zeta(s,t)=\sum_{m=0}^{\infty}(m+t)^{-s}$.
+In this section we consider the polynomial weights $v_{\alpha}(x)=(1+|x|)^{\alpha}$ and derive bounds on $\Phi_{v_{\alpha}}(p)$ in (24).
 Lemma 4.9.
 If $\alpha>\frac{1}{2}$, then
-$\Phi_{v_{\alpha}}(p)\leq p^{-\alpha}\sqrt{2\zeta(2\alpha,\tfrac{1}{2})}\leq p^{-\alpha}2^{\alpha+1/2}\Big{(}1+\frac{1}{4\alpha-2}\Big{)}^{1/2}\lesssim p^{-\alpha}\,.$
-Proof 3.8.
+$\Phi_{v_{\alpha}}(p)\leq p^{-\alpha}2^{\alpha+1/2}\Big(1+\frac{1}{4\alpha-2}\Big)^{1/2}\lesssim p^{-\alpha}\,.$
+Proof 4.10.
 We factor out $p$ to obtain
 $|\Phi_{v_{\alpha}}(p)|^{2}=2\sum_{m=0}^{\infty}\left(1+\left|pm+\tfrac{p}{2}\right|\right)^{-2\alpha}$ $\leq p^{-2\alpha}2\sum_{m=0}^{\infty}\left(\tfrac{1}{p}+m+\tfrac{1}{2}\right)^{-2\alpha}$
-$\leq p^{-2\alpha}2\sum_{m=0}^{\infty}\left(m+\tfrac{1}{2}\right)^{-2\alpha}=p^{-2\alpha}2\zeta(2\alpha,\tfrac{1}{2})\,.$
-The integral test for convergence of series provides, for $\alpha>1/2$,
+$\leq p^{-2\alpha}2\sum_{m=0}^{\infty}\left(m+\tfrac{1}{2}\right)^{-2\alpha}\,.$
+This sum has the shape of the Hurwitz zeta function $\zeta(s,t)=\sum_{m=0}^{\infty}(m+t)^{-s}$ at $t=1/2$. The integral test for convergence of series provides, for $\alpha>1/2$,
 $\zeta(2\alpha,\tfrac{1}{2})\leq 2^{2\alpha}+\int_{0}^{\infty}(x+\tfrac{1}{2})^{-2\alpha}\mathrm{d}x=2^{2\alpha}(1+\tfrac{1}{4\alpha-2})\,.$
 For polynomial weights Theorem 4.7 takes the following shape.
-Corollary 3.9 (polynomial weights).
+Theorem 4.11 (polynomial weights).
 Assume that $f\in W(\mathcal{C},\ell^{2}_{v_{\alpha}})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{v_{\beta}})$ for $\alpha,\beta>\frac{1}{2}$, and $0<h\leq 1\leq p$.
@@
     If the step size satisfies $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$, then
-    $E^{[n]}_{h}(f)\lesssim h^{\beta}\big{(}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\big{)}\,.$
-    Expressed with the number of samples, the error is $E^{[n]}_{h}(f)\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}$.
-Proof 3.10.
+    $E^{[n]}_{h}(f)\lesssim h^{\beta}\big(\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\big)\,$,
+    where the constant in $\lesssim$ is independent of $h,n,p,f$, but may depend on $\alpha,\beta$. Expressed with the number of samples, the error is
+    $E^{[n]}_{h}(f)\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}$
+    with a constant independent of $h,n,p$.
+Part (ii) can be made more explicit by collecting all constants, leading to the ideal choice of $h$ in relation to properties of $f$.
+Corollary 4.12 (explicit constants).
+Under the assumptions of Theorem 4.11, we make the choice
+$h=\left(\frac{c_{\alpha}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}}{c_{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}}\right)^{\frac{1}{\alpha+\beta}}n^{-\frac{\alpha}{\alpha+\beta}}\,$,
+where $c_{s}=2^{2s+1}\left(1+\tfrac{1}{4s-2}\right)^{\frac{1}{2}}$. Then the resulting error bound becomes
+$E^{[n]}_{h}(f)\leq 2\left(c_{\alpha}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}\right)^{\frac{\beta}{\alpha+\beta}}\left(c_{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\right)^{\frac{\alpha}{\alpha+\beta}}n^{-\frac{\alpha\beta}{\alpha+\beta}}\,.$
+[…]
+Proof 4.13.
 (i) The assumption $h\leq 1\leq p$ implies that $\sqrt{1+h}\leq\sqrt{2}$ and $\sqrt{1+p^{-1}}\leq\sqrt{2}$.
-By definition $v_{\alpha}(1)=2^{\alpha}$.
-Using Lemma 3.7, the constant $v_{\alpha}(1)\Phi_{v_{\alpha}}(p)\sqrt{1+h}$ in Theorem 3.5 is then bounded by $2^{2\alpha+1}(1+1/(4\alpha-2))^{1/2}$.
-Likewise for the constant involving $w=v_{\beta}$.
+We note $v_{\alpha}(1)=2^{\alpha}$.
+Using Lemma 4.9, the constant $v_{\alpha}(1)\Phi_{v_{\alpha}}(p)\sqrt{1+h}$ in Theorem 4.7 is then bounded by $2^{2\alpha+1}(1+1/(4\alpha-2))^{1/2}$, likewise for the constant involving $w=v_{\beta}$.
 The error estimate now follows from Theorem 4.7.
@@
 Since $h=\frac{p}{n}$, the choice $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$ leads to $h^{\beta}\asymp p^{-\alpha}$ and the overall error is of order $h^{\beta}$.
-Corollary 3.9 answers the questions (Q1) and (Q2) raised in the introduction whenever $f$ decays polynomially in time and in frequency.
+Theorem 4.11 and Corollary 4.12 answer the questions (Q1) and (Q2) raised in the introduction whenever $f$ decays polynomially in time and in frequency.
 Remark 4.14.
 In the literature, e.g.
-[16], one can find the choice $h\asymp\frac{1}{\sqrt{n}}$.
-It is optimal only for $\alpha=\beta$.
+[4, 20], one can find the choice $h\asymp\frac{1}{\sqrt{n}}$.
+It is optimal for $\alpha=\beta$.
 If $\alpha\neq\beta$, then $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$ leads to better bounds.
-Our analysis reveals that the spacing should depend on the decay of $f$ and $\hat{f}$, and Corollary 3.9 explains how to choose the optimal parameters.
+That is, knowledge on the decay of $f$ and $\hat{f}$ allows us to refine the spacing, and Theorem 4.11 explains how to choose the optimal parameters.
 In our experience, amalgam spaces are the natural theoretical framework when dealing with sampling and periodization.
 Yet one may wish to have conditions that are easier to check in practice.
-One such condition is a pure decay condition $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ as in Theorem 1.1.
-Proof 3.12 (Proof of Theorem 1.1).
-Assume that $|f(x)|\lesssim(1+|x|)^{-a}$ and $\alpha<a-\frac{1}{2}$. We show that $f\in W(\mathcal{C},\ell^{2}_{v_{\alpha}})$. Hölder’s inequality yields
-$\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{2}(1+|l|)^{2\alpha}\lesssim\sup_{x\in\mathbb{R}}|f(x)|^{2}(1+|x|)^{2a}\sum_{l\in\mathbb{Z}}(1+|l|)^{-2(a-\alpha)}\,.$ (21)
+One such condition is a pure decay condition of the form $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ as in Theorem 2.1.
+Proof 4.15 (Proof of Theorem 2.1).
+Assume that $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{a}<\infty$ and $\alpha<a-\frac{1}{2}$. We show that $f\in W(\mathcal{C},\ell^{2}_{v_{\alpha}})$. Hölder’s inequality yields
+(29) $\sum_{l\in\mathbb{Z}}\sup_{x\in[0,1]}\left|f(x+l)\right|^{2}(1+|l|)^{2\alpha}\lesssim\sup_{x\in\mathbb{R}}|f(x)|^{2}(1+|x|)^{2a}\sum_{l\in\mathbb{Z}}(1+|l|)^{-2(a-\alpha)}\,.$
 Since $a-\alpha>\frac{1}{2}$, the series $\sum_{l\in\mathbb{Z}}(1+|l|)^{-2(a-\alpha)}<\infty$ converges, so that
@@
 4.6 Sub-exponential weights
-Next, we consider sub-exponential weights of the form $v_{r,\alpha}(x)=e^{r|x|^{\alpha}}$ as defined in Section 3.2. We first find a bound for $\Phi_{v_{r,\alpha}}(p)$ in (16). We directly derive
-$|\Phi_{v_{r,\alpha}}(p)|^{2}=2\sum_{m=0}^{\infty}e^{-2r\left(pm+\frac{p}{2}\right)^{\alpha}}=e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,.$ (22)
+Next, we consider sub-exponential weights of the form $v_{r,\alpha}(x)=e^{r|x|^{\alpha}}$. We first find a bound for $\Phi_{v_{r,\alpha}}(p)$ in (24). We directly derive
+(30) $|\Phi_{v_{r,\alpha}}(p)|^{2}=2\sum_{m=0}^{\infty}e^{-2r\left(pm+\frac{p}{2}\right)^{\alpha}}=e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,.$
 For fixed $0<\alpha\leq 1$, the series $\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}$ converges for every $p>0$.
 As a function of $p$, it is monotonically decreasing, so that
-$|\Phi_{v_{r,\alpha}}(p)|^{2}\leq e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{1}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,,$ (23)
+(31) $|\Phi_{v_{r,\alpha}}(p)|^{2}\leq e^{-2r(\frac{p}{2})^{\alpha}}2\sum_{m=0}^{\infty}e^{-2r(\frac{1}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}\,$,
 for $p\geq 1$. To derive more explicit bounds, we use the incomplete Gamma function $\Gamma(s,u)=\int_{u}^{\infty}t^{s-1}e^{-t}\mathrm{d}t$.
@@
     $\Phi_{v_{r,\alpha}}(p)\leq e^{-r(\frac{p}{2})^{\alpha}}\Big(2+\frac{2^{\alpha}}{2\alpha^{2}rp^{\alpha}}\Big)^{1/2}\leq e^{-r(\frac{p}{2})^{\alpha}}\Big(2+\frac{1}{\alpha}\Big)^{1/2}\,.$
-In particular, the factor $\big{(}2+\frac{2^{\alpha}}{2\alpha^{2}rp^{\alpha}}\big{)}^{1/2}$ tends to $\sqrt{2}$ when $p\rightarrow\infty$.
-Proof 3.14.
+Note that the factor $\big(2+\frac{2^{\alpha}}{2\alpha^{2}rp^{\alpha}}\big)^{1/2}$ tends to $\sqrt{2}$ when $p\rightarrow\infty$.
+Proof 4.17.
 According to (30), we must bound the series $\sum_{m=0}^{\infty}e^{-2r(\frac{p}{2})^{\alpha}\left((2m+1)^{\alpha}-1\right)}$.
@@
 $\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}$ $\leq 1+\int_{0}^{\infty}e^{-u\left((2x+1)^{\alpha}-1\right)}\mathrm{d}x=1+\frac{1}{2}e^{u}\int_{1}^{\infty}e^{-ux^{\alpha}}\mathrm{d}x\,.$
-A direct check reveals that $-\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},x^{\alpha}u)$ is a primitive of $x\mapsto e^{-ux^{\alpha}}$ and therefore $\int_{1}^{\infty}e^{-ux^{\alpha}}\,\mathrm{d}x=-\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},u)$.
+A direct check reveals that $\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},x^{\alpha}u)$ is a primitive of $x\mapsto e^{-ux^{\alpha}}$ and therefore $\int_{1}^{\infty}e^{-ux^{\alpha}}\,\mathrm{d}x=\frac{1}{\alpha}u^{-\frac{1}{\alpha}}\Gamma(\frac{1}{\alpha},u)$.
 Hence, we derive
-$\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}\leq 1+\frac{1}{2\alpha}\frac{e^{u}}{u^{1/\alpha}}\Gamma(\tfrac{1}{\alpha},u)\,.$ (24)
+(32) $\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}\leq 1+\frac{1}{2\alpha}\frac{e^{u}}{u^{1/\alpha}}\Gamma(\tfrac{1}{\alpha},u)\,.$
 The condition $p\geq\big(\frac{2^{\alpha}}{2r\alpha}\big)^{1/\alpha}$ implies $u=2r(\frac{p}{2})^{\alpha}\geq\frac{1}{\alpha}\geq 1$. In that case, the incomplete Gamma function is bounded by
-$\Gamma(\tfrac{1}{\alpha},u)\leq\frac{1}{\alpha}u^{\frac{1}{\alpha}-1}e^{-u}\,,$ (25)
-see [18, Proposition 2.7 and Page 1275]. For completeness we reproduce the elementary argument:
-Set $a=\alpha^{-1}$. The function $h(t)=(a-1)\ln t-t$ is strictly concave, so it is majorized by its tangent $h_{u}(t)=h(u)+h^{\prime}(u)(u-x)$. Therefore
-[…]
+(33) $\Gamma(\tfrac{1}{\alpha},u)\leq\frac{1}{\alpha}u^{\frac{1}{\alpha}-1}e^{-u}\,$,
+see [23, Proposition 2.7 and Page 1275].
+For completeness we reproduce the elementary argument: set $a=\alpha^{-1}$.
+The function $h(t)=(a-1)\ln t-t$ is strictly concave, so it is majorized by its tangent $h_{u}(t)=h(u)+h^{\prime}(u)(t-u)$ at $u$.
+Therefore
+$\Gamma(a,u)=\int_{u}^{\infty}e^{h(t)}\,dt<\int_{u}^{\infty}e^{h_{u}(t)}\,dt=\frac{u^{a-1}e^{-u}}{1-(a-1)/u}\leq a{u^{a-1}e^{-u}}\,$,
 where the last inequality is due to $(1-(a-1)/u)^{-1}\leq a$ for $u\geq a\geq 1$. This yields (33).
-We substitute (25) into (24), use the assumption $u\leq 1/\alpha$, and obtain
+We substitute (33) into (32), use the assumption $u\geq 1/\alpha$, and obtain
 $\sum_{m=0}^{\infty}e^{-u\left((2m+1)^{\alpha}-1\right)}$ $\leq 1+\frac{1}{2\alpha^{2}u}\leq 1+\frac{1}{2\alpha}\,.$
@@
     $E^{[n]}_{h}(f)\lesssim e^{-s(2h)^{-\beta}}\big(\|f\|_{W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{s,\beta}})}\big)\,.$
-Proof 3.16.
+Proof 4.19.
 The proof is analogous to the one for polynomial weights (Corollary 4.11). We only check the constants. According to Theorem 4.7 and Lemma 4.16, the constant $c_{r,\alpha}$ can be chosen as
@@
 where we used $\sqrt{1+h}\leq\sqrt{2}$. Likewise, we get the specification of $c_{s,\beta}$.
-Proof 3.17 (Proof of Theorem 1.2).
-Assume that $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $r^{\prime}=r-\epsilon<r$.
-Then $f\in W(\mathcal{C},\ell^{1}_{r-\epsilon,\alpha})$ for every $\epsilon>0$ with the same argument as in (21).
-Therefore the theorem is a direct consequence of Corollary 3.15 and this embedding.
+Proof 4.20 (Proof of Theorem 2.3).
+Assume that $\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}<\infty$ and $r^{\prime}=r-\epsilon<r$.
+Then $f\in W(\mathcal{C},\ell^{2}_{r-\epsilon,\alpha})$ for every $\epsilon>0$ with the same argument as in (29).
+Therefore the theorem is a consequence of Corollary 4.18.
 4.7 Mixed weights
@@
 Results for sub-exponential decay of $\hat{f}$ and polynomial decay of $f$ are obtained by switching the roles of $f$ and $\hat{f}$ in the results below.
-Assume that $f\in W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{v_{\beta}})$. We feed the estimates of Lemma 3.7 and 3.13 into Theorem 3.5 and obtain
-$E^{[n]}_{h}(f)\leq e^{r}\big{(}4+\tfrac{2}{\alpha}\big{)}^{1/2}e^{-r(\frac{p}{2})^{\alpha}}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})}+2^{2\beta+1}\big{(}1+\tfrac{1}{4\beta-2}\big{)}^{\frac{1}{2}}h^{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}$ (26)
+Assume that $f\in W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{v_{\beta}})$. We feed the estimates of Lemma 4.9 and 4.16 into Theorem 4.7 and obtain
+(34) $E^{[n]}_{h}(f)\leq e^{r}\big(4+\tfrac{2}{\alpha}\big)^{1/2}e^{-r(\frac{p}{2})^{\alpha}}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{r,\alpha}})}+2^{2\beta+1}\big(1+\tfrac{1}{4\beta-2}\big)^{1/2}h^{\beta}\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}$
 provided that $p\geq\big(\frac{2^{\alpha}}{2r\alpha}\big)^{1/\alpha}$.
@@
 $h=n^{-1}2(\tfrac{\beta}{\alpha r})^{1/\alpha}\mathcal{W}^{1/\alpha}(\tfrac{\alpha n^{\alpha}r}{2^{\alpha}\beta})\,.$
-As for direct decay conditions, if $f$ and $\hat{f}$ are continuous and satisfy $|f(x)|\lesssim e^{-r|x|^{\alpha}}$ and $|\hat{f}(\xi)|\lesssim(1+|\xi|)^{-b}$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$,
-$E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,.$ (27)
+As for direct decay conditions, if $f$ and $\hat{f}$ are continuous and satisfy
+$\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|^{\alpha}}<\infty$ and $\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}<\infty$ with $b>1$, then, for all $r^{\prime}<r$ and $\beta<b-\frac{1}{2}$,
+(35) $E^{[n]}_{h}(f)\lesssim e^{-r^{\prime}(\frac{p}{2})^{\alpha}}\sup_{x\in\mathbb{R}}|f(x)|e^{r|x|}+h^{\beta}\sup_{\xi\in\mathbb{R}}|\hat{f}(\xi)|(1+|\xi|)^{b}\,.$
 This yields Theorem 2.4 in the introduction.
-For the exponential case $\alpha=1$, a related bound for the sup-norm of $\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}$ is derived by Briggs and Henson [4, Theorem 6.6].
-However, they impose extremely restrictive regularity requirements.
-For a comparable bound the periodization $\mathcal{P}_{p}(f{\bf 1}_{[-\frac{p}{2},\frac{p}{2})]})$ must be $b-2$-times continuously differentiable on $\mathbb{R}$.
-This holds only for those interval lengths $p$, where the derivatives at ±$\pm\frac{p}{2}$ coincide.
+For the exponential case $\alpha=1$, a related bound for the sup-norm of $\hat{f}_{\frac{1}{p},n}-\operatorname{\mathcal{F}}\!f_{h,n}$ is derived by Briggs and Henson [5, Theorem 6.6].
+However, they impose restrictive boundary conditions.
+For a comparable bound the periodization $\mathcal{P}_{p}(f{\bf 1}_{[-\frac{p}{2},\frac{p}{2})]})$ must be b2b-2-times continuously differentiable on $\mathbb{R}$.
+These are already violated when $f(p/2)\neq f(-p/2)$, in which case (35) holds only with $\beta=1$ regardless of the global smoothness of $f$.
 5 Interpolation
 So far, we have approximated the Fourier samples $\hat{f}(\tfrac{k}{p})$ by the discrete Fourier transform $h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}kj/n}=\sqrt{p}\,(\mathcal{F}f_{h,n})(k)$, for $k\in[n]$.
-In this section we change the point of view: we now interpolate the vector $\mathcal{F}f_{h,n}$ and study how the interpolating function approximates $\hat{f}$ on whole real line $\mathbb{R}$.
-To do so, we use three exemplary interpolation schemes that are often used in approximation theory.
-The standard procedure starts with a cardinal interpolating function: this is a nice function $\phi$ on $\mathbb{R}$ that satisfies the interpolation condition $\phi(k)=\delta_{k}$.
+In this section we change the point of view: we now interpolate the vector $\sqrt{p}\,\mathcal{F}f_{h,n}$ and study how the interpolating function approximates $\hat{f}$ on whole real line $\mathbb{R}$.
+To do so, we use three exemplary interpolation schemes that are often used in approximation theory, namely piecewise constant and piecewise linear interpolation, and interpolation using sinc functions.
+The standard procedure starts with a cardinal interpolating function: this is a nice function $\phi$ on $\mathbb{R}$ that satisfies the interpolation condition $\phi(k)=\delta_{k,0}$.
 The approximation of $\hat{f}$ on $\mathbb{R}$ from the discrete Fourier transform is then given by
-$\Phi_{h,n}(\xi)=\sqrt{p}\,\sum_{k\in[n]}\mathcal{F}f_{h,n}(k)\phi(p\xi-k)=h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\phi(p\xi-k)\,.$ (28)
-This expression is an approximation of $\hat{f}$ on $\mathbb{R}$ that satisfies $\Phi_{h,n}(\tfrac{k}{n})=\sqrt{p}\,\mathcal{F}f_{h,n}(k)$, and we seek to estimate the global error $\|\hat{f}-\Phi_{h,n}\|_{2}$.
+(36) $\Psi_{h,n}(\xi)=\sqrt{p}\,\sum_{k\in[n]}\mathcal{F}f_{h,n}(k)\phi(p\xi-k)=h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\phi(p\xi-k)\,.$
+This expression is an approximation of $\hat{f}$ on $\mathbb{R}$ that satisfies $\Psi_{h,n}(\tfrac{k}{p})=\sqrt{p}\,\mathcal{F}f_{h,n}(k)$, and we seek to estimate the global error $\|\hat{f}-\Psi_{h,n}\|_{L^{2}}$.
 As the cardinal interpolating function $\phi$ we will use the first two B-splines $B_{1}={\bf 1}_{[-\frac{1}{2},\frac{1}{2})}$ and $B_{2}(x)=(1-|x|){\bf 1}_{[-1,1]}(x)$ and the cardinal sine function $\operatorname*{sinc}(x)=\tfrac{\sin\pi x}{\pi x}$.
-Clearly, these functions satisfy $\phi(k)=\delta_{k}$.
+Clearly, these functions satisfy $\phi(k)=\delta_{k,0}$.
 To keep the notion compact, we set $B_{0}:=\operatorname*{sinc}$.
-For $\phi=B_{1}$, then right-hand side of (28) is a step function, for $\phi=B_{2}$ we obtain a continuous, piecewise linear function (this is how many software would plot the approximating vector $\mathcal{F}f_{h,n}$), and for $\phi=\operatorname*{sinc}$ we obtain a smooth function whose Fourier transform has support in $[-1/2,1/2]$, which is commonly referred to as a bandlimited function.
-4.1 Convergence rates
+For $\phi=B_{1}$, the right-hand side of (36) is a step function, for $\phi=B_{2}$ we obtain a continuous, piecewise linear function (this is how many software programs would plot the approximating vector $\mathcal{F}f_{h,n}$), and for $\phi=\operatorname*{sinc}$ we obtain a smooth function whose Fourier transform has support in [1/2,1/2][-1/2,1/2], which is commonly referred to as a bandlimited function in time-frequency analysis.
 We first derive an error bound for the bandlimited approximation (36) with $\phi=\operatorname*{sinc}$ for general weights.
-This is then specified to the polynomial and the sub-exponential weights, resulting in explicit error rates.
-The approximation rates for the $B$-splines (28) with $i=1,2$ are restricted to $\alpha\leq 1$ and $\alpha\leq 2$, respectively.
+This is then specialized to the polynomial and the sub-exponential weights, resulting in explicit error rates.
+The approximation rates for the $B$-splines (36) with $i=1,2$ are restricted to decay rates $\alpha\leq 1$ and $\alpha\leq 2$, respectively.
 Theorem 5.1.
-Assume that $0<h\leq 1\leq p$ and that the weights $v,w$ satisfy (3.2). If $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$, then
-$\Big{\|}\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$ (29)
-If $\operatorname*{sinc}$ is replaced with $B_{i}$, for $i=1,2$, then the estimate (29) bounds holds for the polynomial weights $v=v_{\alpha}$ subject to the restriction $1/2<\alpha\leq i$.
+Assume that $0<h\leq 1\leq p$ and that the weights $v,w$ satisfy (22). If $f\in W(\mathcal{C},\ell^{2}_{v})$ and $\hat{f}\in W(\mathcal{C},\ell^{2}_{w})$, then
+(37) $\hskip-4.30554pt\Big\|\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\xi-k)\Big\|_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}+\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
+If $\operatorname*{sinc}$ is replaced with $B_{i}$, for $i=1,2$, then the estimate (37) still holds for the polynomial weight $v=v_{\alpha}$ subject to the restriction $1/2<\alpha\leq i$.
 We note that the error is of the same form as in Theorem 4.7.
-In particular, for polynomial weights $v=v_{\alpha}$ and $w=v_{\beta}$ with $\alpha,\beta>\frac{1}{2}$ we obtain $\Phi_{v}(p)\lesssim p^{-\alpha}$ and $\Phi_{w}(h^{-1})\lesssim h^{\beta}$ (see Lemma 3.7).
-For sub-exponential weights $v=v_{r,\alpha}$ and $w=v_{s,\beta}$ with $0<\alpha,\beta\leq 1$ and $r,s>0$ we obtain $\Phi_{v}(p)\lesssim e^{-r(\frac{p}{2})^{\alpha}}$ and $\Phi_{w}(h^{-1})\lesssim e^{-s(\frac{1}{2h})^{\beta}}$ (see Lemma 3.13).
-Proof 4.2.
+In particular, (i) for the polynomial weights $v=v_{\alpha}$ and $w=v_{\beta}$ with $\alpha,\beta>\frac{1}{2}$ we obtain $\Phi_{v}(p)\lesssim p^{-\alpha}$ and $\Phi_{w}(h^{-1})\lesssim h^{\beta}$ (see Lemma 4.9), and the choice $h\asymp n^{-\frac{\alpha}{\alpha+\beta}}$ leads to
+$\Big\|\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\xi-k)\Big\|_{L^{2}}\lesssim n^{-\frac{\alpha\beta}{\alpha+\beta}}\left(\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}+\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{v_{\beta}})}\right)\,.$
+(ii) For sub-exponential weights $v=v_{r,\alpha}$ and $w=v_{s,\beta}$ with $0<\alpha,\beta\leq 1$ and $r,s>0$ we obtain $\Phi_{v}(p)\lesssim e^{-r(\frac{p}{2})^{\alpha}}$ and $\Phi_{w}(h^{-1})\lesssim e^{-s(\frac{1}{2h})^{\beta}}$ (see Lemma 4.16).
+(iii) For exponential weights $v(x)=e^{r|x|},w(x)=e^{s|x|}$ and the choice $p=\sqrt{n},h=1/\sqrt{n}$ we obtain the root-exponential convergence
+$\Big\|\hat{f}-h\!\!\sum_{k,j\in[n]}\!\!f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\operatorname*{sinc}(p\xi-k)\Big\|_{L^{2}}\lesssim e^{-d\sqrt{n}}\,$,
+as in  [28].
+Proof 5.2 (Proof of Theorem 5.1).
 The proof is based on the following three approximation steps:
-$\hat{f}\,\overset{(s1)}{\approx}\,\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)$ $\,\overset{(s2)}{\approx}\,\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)$
-$\,\overset{(s3)}{\approx}\,h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\cdot-k)\,.$ (30)
-We derive bounds for the steps (s2) and (s3) with respect to general weights simultaneously for $i=0,1,2$.
-(s3) For $i=0,1$, the system $\{\sqrt{p}B_{i}(p\cdot-k)\}_{k\in\mathbb{Z}}$ is orthonormal in $L^{2}(\mathbb{R})$ and it is still a Riesz sequence for $i=2$ with Riesz bounds independent of $p$ (this can be verified by direct computation).
+$\hat{f}(\xi)\,\overset{(s1)}{\approx}\,\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$ $\,\overset{(s2)}{\approx}\,\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$
+(38) $\,\overset{(s3)}{\approx}\,h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\,.$
+Here (s1) is the interpolation error, (s2) the truncation error, and (s3) the discretization error.
+We derive bounds for the error in the steps (s2) and (s3) with respect to general weights simultaneously for $i=0,1,2$.
+(s3) For $i=0,1$, the system $\{\sqrt{p}B_{i}(p\xi-k)\}_{k\in\mathbb{Z}}$ is orthonormal in $L^{2}(\mathbb{R})$ and it is still a Riesz sequence for $i=2$ with Riesz bounds independent of $p$ (this can be verified by direct computation).
 Therefore, we conclude that
-$\Big{\|}\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\cdot-k)\Big{\|}_{L^{2}}$
-$=\|\sum_{k\in[n]}\frac{1}{\sqrt{p}}\Big{(}\hat{f}(\frac{k}{p})-\mathcal{F}f_{h,n}(k)\Big{)}\,\sqrt{p}\,B_{i}(p\cdot-k)\|_{L}^{2}\asymp\big{\|}\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\big{\|}\ =E^{[n]}_{h}(f)\,,$
+$\Big\|\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\Big\|_{L^{2}}$
+$=\Big\|\sum_{k\in[n]}\Big(\tfrac{1}{\sqrt{p}}\hat{f}(\tfrac{k}{p})-\mathcal{F}f_{h,n}(k)\Big)\,\sqrt{p}\,B_{i}(p\xi-k)\Big\|_{L^{2}}\asymp\big\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\big\|\ =E^{[n]}_{h}(f)\,$,
 and this is just the discrete approximation error that has been treated in Theorem 4.7.
 (s2) Using orthonormality or the Riesz basis property again, we observe
-$\Big{\|}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}^{2}_{L^{2}}\asymp p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}\,.$
+$\Big\|\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big\|^{2}_{L^{2}}\asymp p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}\,.$
 The monotonicity (22b) of the weight $w$ implies
 $p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}\lesssim p^{-1}|w(\tfrac{n}{2p})|^{-2}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}|w(\tfrac{k}{p})|^{2}\,.$
-The sum on the right-hand side is bounded as in estimates at the end of the proof of Lemma 3.3. We deduce
+The sum on the right-hand side is bounded as in the estimates at the end of the proof of Lemma 4.5. We deduce
 $p^{-1}\sum_{k\in\mathbb{Z}\setminus[n]}|\hat{f}(\tfrac{k}{p})|^{2}|w(\tfrac{k}{p})|^{2}\leq p^{-1}\lceil p\rceil|w(1)|^{2}\|\hat{f}\|^{2}_{W(\mathcal{C},\ell^{2}_{w})}\,.$
 As $|w(\tfrac{n}{2p})|^{-1}=|w(\tfrac{1}{2h})|^{-1}\leq\big(2\sum_{m=0}^{\infty}|w(\frac{m}{h}+\frac{1}{2h})|^{-2}\big)^{1/2}=\Phi_{w}(h^{-1})$ we obtain
-$\Big{\|}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
-(s1) The error $\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\|_{L^{2}}$ relates to classical interpolation of $\hat{f}$ from its samples [5, 6].
-Since we consider general weights only for $B_{0}=\operatorname*{sinc}$, we carry out the calculations for $B_{0}$.
+$\Big\|\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big\|_{L^{2}}\lesssim\Phi_{w}(h^{-1})\|\hat{f}\|_{W(\mathcal{C},\ell^{2}_{w})}\,.$
+(s1) The interpolation error $\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\|_{L^{2}}$ relates to classical interpolation of $\hat{f}$ from its samples [6, 7, 28].
+We carry out the calculations for $B_{0}=\operatorname*{sinc}$.
 Direct computations and Plancherel’s formula lead to
-$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\cdot-k)\Big{\|}^{2}_{L^{2}}$ $=\Big{\|}f-\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}\,\cdot\,\frac{k}{p}}{\bf 1}_{[-\frac{p}{2},\frac{p}{2}]}\Big{\|}^{2}_{L^{2}}$
+$\Big\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\xi-k)\Big\|^{2}_{L^{2}}$ $=\Big\|f-\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}x\frac{k}{p}}{\bf 1}_{[-\frac{p}{2},\frac{p}{2}]}\Big\|^{2}_{L^{2}}$
 $=\int_{|x|>\frac{p}{2}}\!|f(x)|^{2}\mathrm{d}x+\!\int_{-\frac{p}{2}}^{\frac{p}{2}}\!\bigg|f(x)-\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}x\frac{k}{p}}\bigg|^{2}\mathrm{d}x\,.$
@@
 $\int_{-\frac{p}{2}}^{\frac{p}{2}}\Big|f(x)-p^{-1}\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})e^{2\pi\mathrm{i}x\frac{k}{p}}\Big|^{2}\mathrm{d}x$ $=\|f-\mathcal{P}\!_{p}f\|^{2}_{L^{2}(-\frac{p}{2},\frac{p}{2})}\lesssim\Phi^{2}_{v}(p)\|f\|^{2}_{W(\mathcal{C},\ell^{2}_{v})}\,.$
-This provides the required bound
-$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$
-To treat the case $i=1,2$ of piecewise constant and piecewise linear interpolation for the polynomial weights $v=v_{\alpha}$ and $w=v_{\beta}$, we use results from [15, Thm 3 and Lemma 3].
+This provides the bound for the interpolation error as
+$\Big\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\operatorname*{sinc}(p\xi-k)\Big\|_{L^{2}}\lesssim\Phi_{v}(p)\|f\|_{W(\mathcal{C},\ell^{2}_{v})}\,.$
+To treat the cases $i=1,2$ of piecewise constant and piecewise linear interpolation for the polynomial weight $v=v_{\alpha}$, we use results from [19, Theorem 3 and Lemma 3].
 For $\alpha\leq i,i=1,2,$ these imply
-$\Big{\|}\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\Big{\|}_{L^{2}}\lesssim p^{-\alpha}\|f\|_{L^{2}_{v_{\alpha}}}\,.$ (31)
-The continuous embedding $W(\mathcal{C},\ell^{2}_{v_{i}})\subseteq L^{2}_{v_{i}}$ yields $\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\cdot-k)\|_{L^{2}}\lesssim p^{-i}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{i}})}$.
+(39) $\Big\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\Big\|_{L^{2}}\lesssim p^{-\alpha}\|f\|_{L^{2}_{v_{\alpha}}}\,.$
+The continuous embedding $W(\mathcal{C},\ell^{2}_{v_{\alpha}})\subseteq L^{2}_{v_{\alpha}}$ yields
+$\|\hat{f}-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)\|_{L^{2}}\lesssim p^{-\alpha}\|f\|_{W(\mathcal{C},\ell^{2}_{v_{\alpha}})}\,.$
 The final error is the sum of the three errors (s1), (s2), and (s3).
 Remark 5.3.
-The restriction to $\alpha=i$, for $i=1,2$, in Part (b) of Theorem 4.1 stems from the bound (31).
+The restriction to $\alpha\leq i$, for $i=1,2$, in the second part of Theorem 5.1 stems from the bound (39).
 For stronger polynomial decay $\alpha$, the splines $B_{1}$ and $B_{2}$ need to be replaced by a suitable function $\phi_{\alpha}$, see [6] for examples.
 We need at least that $\phi_{\alpha}$ interpolates on $\mathbb{Z}$ and that its integer shifts form a Riesz-sequence in $L^{2}$.
-The proof for the interpolating function $\phi=\operatorname*{sinc}$ works for all polynomials weights $v_{\alpha}$ with $\alpha>\frac{1}{2}$.
-Although the $\operatorname*{sinc}$-function has a bad reputation in numerical analysis due to its poor decay, it works well in the context of Fourier approximation.
-In particular, since we use only finite sums, the slow decay of $\operatorname*{sinc}$ does not cause any convergence issues.
-We refer to [21, 22] for $\operatorname*{sinc}$-methods in numerical analysis.
-5 Minimal requirements
-In this section we investigate under which conditions the error $E^{[n]}_{h}(f)$ tends to zero.
-Although this question may not be of practical value, it is of fundamental interest to know for which function class the Fourier transform can be successfully approximated by the discrete Fourier transform.
-We answer this question for the norm $\sqrt{p}\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}$ representing the pointwise error.
+The proof for the interpolating function $\phi=\operatorname*{sinc}$ holds for all polynomial weights $v_{\alpha}$ with $\alpha>\frac{1}{2}$.
+The $\operatorname*{sinc}$-function arises naturally in Fourier approximation, but its slow decay poses challenges in numerical analysis.
+Nevertheless, it has been successfully employed in various numerical methods – see Stenger’s work [27, 28] for comprehensive treatments of $\operatorname*{sinc}$-based techniques.
+6 Minimal requirements for convergence
+In this section we aim to identify the weakest conditions under which the Fourier transform can be successfully approximated by the discrete Fourier transform.
+Although this question may not be of immediate practical value, it is a fundamental question of intrinsic mathematical interest.
+To identify weak conditions, we now choose the weaker norm
+(40) $\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\,$
+and ask for sufficient conditions on $f$ and $\hat{f}$ such that (40) converges to 0 when $h\rightarrow 0$ and $p\rightarrow\infty$.
+It turns out that the same general conditions for the validity of the Poisson summation formula also guarantee the convergence.
 Proposition 6.1.
@@
 $\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\,\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|=0\,.$
-Proof 5.2.
-Following the general approach outlined in Section 3.1 we decompose the error $\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}$ into a time and a frequency component. As in (13), we have
-$\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}\leq\|\hat{f}_{\frac{1}{p},n}-(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\|\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}-\operatorname{\mathcal{F}}\!f_{h,n}\|_{\infty}\,.$ (32)
+The limit above is taken in the sense that $h\to 0$ and $p^{-1}\to 0$ in arbitrary relation to each other. In particular, it implies that $n=\frac{p}{h}\rightarrow\infty$.
+Proof 6.2.
+Following the general approach outlined in Section 4.1, we decompose the error $\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|=\sqrt{p}\,\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}$ into a time and a frequency component.
+Since we are taking limits, we may assume without loss of generality that $h\leq 1$.
+As in (4.6), we have
+(41) $\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}\leq\|\hat{f}_{\frac{1}{p},n}-(\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\|\mathcal{F}(\mathcal{P}\!_{p}f)_{h,n}-\operatorname{\mathcal{F}}\!f_{h,n}\|_{\infty}\,.$
 Since the discrete Fourier transform satisfies $\|\mathcal{F}y\|_{\infty}\leq\frac{1}{\sqrt{n}}\|y\|_{1}$ for $y\in\mathbb{C}^{n}$, we obtain
-$\sqrt{p}\|\hat{f}_{\frac{1}{p},n}-\mathcal{F}f_{h,n}\|_{\infty}\leq\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}\,,$ (33)
-which again splits the error into a time component and a frequency component.
+(42) $\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\leq\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}+\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}\,$,
+which splits the error into a time component and a frequency component as before.
 It remains to obtain suitable decay of samples of $\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f}$ and $f-\mathcal{P}\!_{p}f$.
 We start with the frequency component $\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$.
-The observation $|\frac{k}{p}|\leq\frac{1}{2h}$, for $k\in[n]$, and assuming without loss of generality that $h\leq 1$ lead to
-$\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$ $=\sup_{k\in[n]}\Big{|}\sum_{l\in\mathbb{Z}\setminus\{0\}}\hat{f}(\tfrac{k}{p}+h^{-1}l)\Big{|}\leq\sum_{|l|\geq\frac{1}{2h}}\sup_{\xi\in[0,1]}|\hat{f}(\xi+l)|\,.$ (34)
+The observation $|\frac{k}{p}|\leq\frac{1}{2h}$ for $k\in[n]$ and the assumption $h\leq 1$ lead to
+(43) $\sqrt{p}\|(\hat{f}-\mathcal{P}\!_{h^{-1}}\hat{f})_{\frac{1}{p},n}\|_{\infty}$ $=\sup_{k\in[n]}\Big|\sum_{l\in\mathbb{Z}\setminus\{0\}}\hat{f}(\tfrac{k}{p}+h^{-1}l)\Big|\leq\sum_{|l|\geq\frac{1}{2h}}\sup_{\xi\in[0,1]}|\hat{f}(\xi+l)|\,.$
 Since $\hat{f}\in W(\mathcal{C},\ell^{1})$, this sum tends to zero for $h\rightarrow 0$.
@@
 $\sqrt{h}\|(f-\mathcal{P}\!_{p}f)_{h,n}\|_{1}\leq h\lceil h^{-1}\rceil\sum_{|l|\geq\frac{p}{2}}\sup_{x\in[0,1]}|f(x+l)|\,.$
-Since $f\in W(\mathcal{C},\ell^{1})$, the series tends to zero for $p\rightarrow\infty$.
-The factor $h\lceil h^{-1}\rceil\leq(1+h)$ does not cause any issues because $h$ is small, after all, we let $h\rightarrow 0$.
-Since $\|y\|_{2}\leq\sqrt{n}\,\|y\|_{\infty}$ for $y\in\mathbb{C}^{n}$, Proposition 5.1 also implies the following $\ell^{2}$-error:
-$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sqrt{h}\,E^{[n]}_{h}(f)\leq\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sqrt{h}\frac{\sqrt{n}}{\sqrt{p}}\,\sup_{k\in[n]}\Big{|}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{|}\rightarrow 0\,.$
-for $f,\hat{f}\in W(\mathcal{C},\ell^{1})$.
-In a similar spirit we can ask when the interpolation in (28) with the $B$-splines $B_{1}$ or $B_{2}$ converges to zero.
+Since $f\in W(\mathcal{C},\ell^{1})$, the series tends to zero for $p\rightarrow\infty$. The factor $h\lceil h^{-1}\rceil\leq(1+h)\leq 2$ does not cause any issues because $h\rightarrow 0$.
+In previous sections, we considered the $\ell^{2}$-error $E^{[n]}_{h}(f)$, which is stronger than (40).
+Since $\|y\|_{2}\leq\sqrt{n}\,\|y\|_{\infty}$ for $y\in\mathbb{C}^{n}$, Proposition 6.1 also implies that for $f,\hat{f}\in W(\mathcal{C},\ell^{1})$,
+$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sqrt{h}\,E^{[n]}_{h}(f)\leq\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\rightarrow 0\,$,
+so that an additional factor $\sqrt{h}$ is needed for convergence.
+Parallel to Section 5 we now investigate when the interpolation error of $\hat{f}$ in (36) converges to zero.
 Proposition 6.3.
-If $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then
-$\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\,\sup_{\xi\in\mathbb{R}}\Big{|}\hat{f}(\xi)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\Big{|}=0\,,\qquad i=1,2.$ (35)
-A similar result under a stronger condition on $f$ and $\hat{f}$ and the choice $p=1/h=\sqrt{n}$ was derived in [16] for the piecewise linear interpolation with $B_{2}$.
-Proposition 5.3 improves the main result in [16] by extending the approximation result to a larger class of functions.
-For the rather non-trivial comparison of the condition $f,\hat{f}\in W(\mathcal{C},\ell^{1})$ to the one in [16] we refer to [17, Theorem 2]).
-Proof 5.4.
-As in the proof of Theorem 4.1 we split the approximation error into discretization, truncation, and approximation via the discrete Fourier transform as follows:
+If $f,\hat{f}\in W(\mathcal{C},\ell^{1})$, then for interpolation with step functions $B_{1}$ and with piecewise linear functions $B_{2}$ we obtain
+(44) $\lim_{\begin{subarray}{c}h\rightarrow 0\\ p\rightarrow\infty\end{subarray}}\,\sup_{\xi\in\mathbb{R}}\Big|\hat{f}(\xi)-h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\Big|=0\,,\qquad i=1,2.$
+Proof 6.4.
+As in the proof of Theorem 5.1 we split the approximation error into interpolation, truncation, and discretization via the discrete Fourier transform as follows:
 $\hat{f}(\xi)\,\overset{(a1)}{\approx}\,\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$ $\,\overset{(a2)}{\approx}\,\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})B_{i}(p\xi-k)$
 $\,\overset{(a3)}{\approx}\,h\sum_{k,j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}B_{i}(p\xi-k)\,.$
-To estimate the error in $(a1)$, we use the fact that the integer shifts of $B_{i}$, $i=1,2$, form a partition of unity, so that $\operatorname*{supp}(B_{i})\subseteq[-1,1]$ leads to
+To estimate the interpolation error $(a1)$, we use the fact that $B_{i}$ has compact support, $\operatorname*{supp}(B_{i})\subseteq[-1,1]$, and that the integer shifts of $B_{i}$, $i=1,2$, form a partition of unity.
+This leads to
 $\Big|\hat{f}(\xi)-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\,B_{i}(p\xi-k)\Big|$ $=\Big|\sum_{k\in\mathbb{Z}}(\hat{f}(\xi)-\hat{f}(\tfrac{k}{p}))\,B_{i}(p\xi-k)\Big|$
-$\leq\Big{|}\sum_{\begin{subarray}{c}k\in\mathbb{Z}\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}|\hat{f}(\xi)-\hat{f}(\tfrac{k}{p})|B_{i}(p\xi-k)\Big{|}$
-$\leq\sup_{k\in\mathbb{Z}}\sup_{\xi\in\mathbb{R}:|\xi-\frac{k}{p}|\leq\frac{1}{p}}\Big{(}\sum_{k\in\mathbb{Z}}B_{i}(p\xi-k)\Big{)}$
-$\leq\sup_{\xi,\eta\in\mathbb{R}:|\xi-\eta|<\frac{1}{p}}|\hat{f}(\xi)-f(\eta)|$
+$\leq\sum_{\begin{subarray}{c}k\in\mathbb{Z}\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}|\hat{f}(\xi)-\hat{f}(\tfrac{k}{p})|\,B_{i}(p\xi-k)$
+$\leq\sup_{k\in\mathbb{Z}}\sup_{\xi\in\mathbb{R}:|\xi-\frac{k}{p}|\leq\frac{1}{p}}|\hat{f}(\xi)-\hat{f}(\tfrac{k}{p})|\sum_{k\in\mathbb{Z}}\,B_{i}(p\xi-k)$
+$\leq\sup_{\xi,\eta\in\mathbb{R}:|\xi-\eta|\leq\frac{1}{p}}|\hat{f}(\xi)-\hat{f}(\eta)|\,.$
 Since $f\in W(\mathcal{C},\ell^{1})$ is uniformly continuous and $p\rightarrow\infty$, we conclude that
 $\lim_{p\rightarrow\infty}\sup_{\xi\in\mathbb{R}}\Big|\hat{f}(\xi)-\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\,B_{i}(p\xi-k)\Big|=0\,.$
-For the error in $(a2)$, we use $\|B_{i}\|_{\infty}=1$ and find for all $\xi\in\mathbb{R}$ that
+For the truncation error $(a2)$ we use $\|B_{i}\|_{\infty}=1$ and find for all $\xi\in\mathbb{R}$ that
 $\Big|\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\,B_{i}(p\xi-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})\,B_{i}(p\xi-k)\Big|$ $=\Big|\sum_{k\in\mathbb{Z}\setminus[n]}\hat{f}(\tfrac{k}{p})\,B_{i}(p\xi-k)\Big|$
-$\leq\|B_{i}\|_{\infty}\sum_{\begin{subarray}{c}k\in\mathbb{Z}\setminus[n]\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}\big{|}\hat{f}(\tfrac{k}{p})\big{|}\leq\sum_{k\in\mathbb{Z}\setminus[n]}\big{|}\hat{f}(\tfrac{k}{p})\big{|}\,.$
+$\leq\|B_{i}\|_{\infty}\sum_{\begin{subarray}{c}k\in\mathbb{Z}\setminus[n]\\ |\xi-\frac{k}{p}|\leq\frac{1}{p}\end{subarray}}\big|\hat{f}(\tfrac{k}{p})\big|$
+$\leq\sum_{k\in\mathbb{Z}\setminus[n]}\big|\hat{f}(\tfrac{k}{p})\big|\,.$
 Since $p\geq 1$ and $\frac{|k|}{p}\geq\frac{n}{2}\,\frac{1}{nh}=\frac{1}{2h}$ for $k\in[n]$, the last sum is estimated as in (43) by
@@
 $\lim_{h\rightarrow 0}\sup_{\xi\in\mathbb{R}}\Big|\sum_{k\in\mathbb{Z}}\hat{f}(\tfrac{k}{p})\,B_{i}(p\xi-k)-\sum_{k\in[n]}\hat{f}(\tfrac{k}{p})\,B_{i}(p\xi-k)\Big|=0\,.$
-In order to estimate $(a3)$, we compute
-$\bigg{|}\sum_{k\in[n]}\Big{(}\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big{)}B_{i}(p\xi-k)\bigg{|}$
+For the discretization error $(a3)$, we estimate
+$\hskip-56.9055pt\bigg|\sum_{k\in[n]}\Big(\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big)\,B_{i}(p\xi-k)\bigg|$
 $\qquad\leq\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\,\Big|\sum_{k\in[n]}B_{i}(p\xi-k)\Big|$
 $\qquad\leq\sup_{k\in[n]}\Big|\hat{f}(\tfrac{k}{p})-h\sum_{j\in[n]}f(hj)e^{-2\pi\mathrm{i}\frac{kj}{n}}\Big|\,$,
-were we have used again that the integer translates of the cardinal $B$-splines form a partition of unity.
+where we have used again that the integer translates of the cardinal $B$-splines form a partition of unity.
 According to Proposition 6.1 this term vanishes when $p\rightarrow\infty$ and $h\rightarrow 0$.
 The final error is the sum of the three errors (a1), (a2), and (a3).
-Appendix A Time frequency symmetry of the error functional
-We prove the time-frequency symmetry of the error $E^{[n]}_{h}$.
-Lemma A.1.
-If $p=nh$, then
-$E^{[n]}_{h}(f)=E^{[n]}_{p^{-1}}(\bar{\hat{f}}),\qquad f,\hat{f}\in W(\mathcal{C},\ell^{1})\,.$ (36)
-Hence, the time frequency decomposition of the error matches the time-frequency symmetry of $E^{[n]}_{h}$.
-Proof A.2.
-To verify (36), we recall that $(\bar{\hat{f}})\,\widehat{}=\bar{f}$ and that $\mathcal{F}$ is unitary, so that
-$E^{[n]}_{p^{-1}}(\bar{\hat{f}})$ $=\|(\overline{\widehat{f_{h,n}}})\,\widehat{}-\mathcal{F}\bar{\hat{f}}_{\frac{1}{p},n}\|=\|\bar{f}_{h,n}-\mathcal{F}\bar{\hat{f}}_{\frac{1}{p},n}\|$
-=*$=\|\mathcal{F}^{*}\bar{f}_{h,n}-\bar{\hat{f}}_{\frac{1}{p},n}\|=E^{[n]}_{p^{-1}}(\bar{\hat{f}})\,.$
-The discrete Fourier transform also satisfies $\overline{\mathcal{F}y}=\mathcal{F}^{*}\bar{y}$, where conjugation is meant entry-wise. Conjugation does not affect the norm, so that we obtain (36).
+Remark 6.5.
+(i) The formulation of Proposition 2.6 in the introduction follows, because the condition $\sup_{x\in\mathbb{R}}|f(x)|(1+|x|)^{1+\epsilon}<\infty$ implies that $f\in W(\mathbb{C},\ell^{1})$.
+(ii) Convergence in a stronger norm for the choice $p=1/h=\sqrt{n}$ and under stronger conditions on $f$ and $\hat{f}$ was derived in [20] for the piecewise linear interpolation with $B_{2}$.
+Proposition 6.3 extends the applicability of the main result in [20] to a larger class of functions.
+For the rather non-trivial comparison of the condition $f,\hat{f}\in W(\mathcal{C},\ell^{1})$ to the one in [20] we refer to [22, Theorem 2].
+(iii) For $B_{0}=\operatorname*{sinc}$ the above proof of Proposition 6.3 breaks down at several steps, because the integer shifts of $\operatorname*{sinc}$ are not summable.
+We do not know whether Proposition 6.3 remains true in this case.